### v1

# Affective Computing in the Era of Large Language Models: A Survey from the NLP Perspective

Abstract
Affective Computing (AC), integrating computer science, psychology, and cognitive science knowledge, aims to enable machines to recognize, interpret, and simulate human emotions. To create more value, AC can be applied to diverse scenarios, including social media, finance, healthcare, education, etc. Affective Computing (AC) includes two mainstream tasks, i.e., Affective Understanding (AU) and Affective Generation (AG). Fine-tuning Pre-trained Language Models (PLMs) for AU tasks has succeeded considerably. However, these models lack generalization ability, requiring specialized models for specific tasks. Additionally, traditional PLMs face challenges in AG, particularly in generating diverse and emotionally rich responses. The emergence of Large Language Models (LLMs), such as the ChatGPT series and LLaMA models, brings new opportunities and challenges, catalyzing a paradigm shift in AC. LLMs possess capabilities of in-context learning, common sense reasoning, and advanced sequence generation, which present unprecedented opportunities for AU. To provide a comprehensive overview of AC in the LLMs era from an NLP perspective, we summarize the development of LLMs research in this field, aiming to offer new insights. Specifically, we first summarize the traditional tasks related to AC and introduce the preliminary study based on LLMs. Subsequently, we outline the relevant techniques of popular LLMs to improve AC tasks, including Instruction Tuning and Prompt Engineering. For Instruction Tuning, we discuss full parameter fine-tuning and parameter-efficient methods such as LoRA, P-Tuning, and Prompt Tuning. In Prompt Engineering, we examine Zero-shot, Few-shot, Chain of Thought (CoT), and Agent-based methods for AU and AG. To clearly understand the performance of LLMs on different Affective Computing tasks, we further summarize the existing benchmarks and evaluation methods. Finally, we discuss challenges from an NLP perspective, such as ethical considerations and effective evaluation, and outline future directions. Our survey seeks to provide valuable insights and references for researchers in this rapidly evolving domain.
Index Terms:
Affective Computing, Large Language Models, Instruction Tuning, Prompt Engineering, Benchmarks.
1 Introduction
Affective Computing (AC), encompassing the study and development of systems capable of recognizing, interpreting, processing, and simulating human affects through natural language and other modalities, seeks to bridge the gap between human emotions and machine understanding [1, 2, 3]. The AC task can be broadly categorized into two mainstream tasks: the Affective Understanding (AU) task [4] and the Affective Generation (AG) task [5, 6]. Affective Understanding focuses on recognizing and interpreting human emotions, covering Sentiment Analysis [7, 8], Sarcasm Detection [9, 10], and so on. Affective Generation involves creating responses or content that can elicit specific emotional responses in humans to generate expressions with emotionally nuanced, including Emotional-aware Dialogue Generation[11, 12], Creative Content Generation, etc. With the emergence of Large Language Models (LLMs)[13, 14, 15], the Affective Computing task is experiencing a paradigm shift to explore and model human emotions in unprecedented ways.
Figure 1: Comparison of approaches to AC Tasks in the PLMs era and the LLMs era. In the PLMs era, specific datasets were typically used to fine-tune PLMs. Each specific dataset usually requires separate model fine-tuning to achieve optimal performance. When addressing AC tasks, PLMs often utilized sentiment-aware modules to assist in AU and AG tasks. The deep neural networks (DNN) [16] used for AC tasks could have various structures. For example, a multilayer perceptron (MLP) [17] might be used for AU tasks, while a long short-term memory network (LSTM) [18] might be used for AG tasks. In contrast, LLMs possess strong generalization capabilities, allowing them to handle multiple tasks simultaneously without needing for separate fine-tuning for each specific task. When using adapters for instruction tuning, the parameters of the LLM are typically frozen to reduce resource consumption and improve efficiency.
Prior to the advent of LLMs, the dominant paradigm of Affective Computing involves fine-tuning Pre-trained Language Models (PLMs) [19], such as BERT [20] and RoBERTa [21], on labelled datasets for specific tasks to improve performance. While the above training paradigm has achieved numerous successes in the tasks related to Affective Understanding, it still is inherently limited by the quality and scale of manually annotated data and struggles to generalize to new domains [22]. The rise of LLMs research, as demonstrated in Figure 1, profoundly transforms the research paradigm of AC, expanding task variety and surpassing the constraints of previous studies, which often focus on particular datasets and single-task the performance improvements [23, 24]. This evolution is accompanied by developing novel approaches that exploit the vast knowledge of LLMs [25] to simultaneously boost the performance of various tasks related to Affective Computing.
The auto-regressive nature and massive pre-training of large language models give them an advantage in processing dialogue inputs and generating coherent responses [26, 27]. Various techniques are proposed to fully harness and enhance the capabilities of LLMs, including instruction tuning and prompt engineering. Specifically, Instruction Tuning [28, 29] involves fine-tuning a pre-trained language model on a dataset where the tasks are explicitly defined through instructions. For example, Flan-T5 [28] specific designs instructions for different tasks, e.g., Sentiment Analysis, Commonsense Reasoning, Question Answering, and so on. Instruction Tuning is a form of targeted training that aims to make the model more adept at following complex and varied instructions across different domains. Prompt Engineering [30, 31], on the other hand, is more about crafting the input given to the model to elicit the desired output without modifying the model itself. In-context Learning can further mine a model to learn and apply knowledge based on the instances provided directly within the prompt, without any prior explicit training on these specific tasks [32]. Although LLMs offer new opportunities and show initial success in various tasks, a comprehensive review that consolidates and summarizes the recent advancements in AC with LLMs from an NLP perspective still needs to be present. Existing surveys either focus narrowly on specific tasks like sentiment analysis or a limited set of classic affective computing tasks [23, 33, 34, 35, 36], or they concentrate on traditional, non-LLM-based methods [19, 37].
To mitigate the gap, it is essential to provide a comprehensive overview of how LLMs are being applied across the diverse spectrum of tasks related to AC, along with the unique opportunities and challenges they bring. Accordingly, we investigate the capabilities and limitations of LLMs in AC tasks, covering both Affective Understanding and Generation. While our focus is primarily on NLP, we also include text-centric multimodal affective computing tasks, recognizing the growing importance of integrating multiple modalities in AC. We begin by summarizing early studies that laid the groundwork for understanding the capabilities and limitations of LLMs in AC. Building on this, we explore two critical technologies for adapting LLMs to AC tasks: instruction tuning and prompt engineering. It involves reviewing various fine-tuning methods designed to enhance LLMs’ performance on AC tasks, followed by an exploration of different prompting strategies aimed at eliciting desired emotional responses from LLMs. To deliver a comprehensive assessment, we further summarize benchmarks and evaluation methods for systematically evaluating LLMs’ performance in AC, including the innovative use of LLMs as evaluation tools. Finally, we enumerate open challenges and outline promising future research directions, offering insights into this rapidly advancing field’s continuous evolution and potential.
2 Tasks
The mainstream tasks of Affective Computing (AC) include Affective Understanding (AU) [4] and Affective Generation (AG) [5, 6]. The primary traditional tasks and related subtasks of AU and AG are summarized in Table I.
TABLE I: The traditional tasks of Affective Computing from an NLP perspective, with representative methods in the LLM era.
AC Tasks Subtasks Detailss
AU Sentiment Analysis (SA) Polarity Classification (PC) USA [38], WisdoM [39], RP-CoT [40], Sun et.al [41]
Emotional Classification (EC) Zhang et.al [35], Stigall et.al [42]
Aspected-Based Sentiment Analysis (ABSA) InstructABSA [43], SCRAP [44]
Emotional Intensity Detection (EID) Amin et.al [45]
Implicit Sentiment Analysis (ISA) THOR [46]
Emotion Recognition in Conversation (ERC) UniMSE [47], DialogueLLMzhang2023dialoguellm, CKERC [48], InstructERC [49]
Emotion Cause Paired Extract (ECPE) DECC [50], ECR-Chain [51]
Subjective Text Analysis (STA) Suicide Tendency Detection (STD) Zhang et.al [35], Amin et.al [45]
Personality Assessment (PA) Zhang et.al [35], Amin et.al [45]
Toxicity Detection (TD) Zhang et.al [35], Amin et.al [45]
Sarcasm Detection (SD) Hoffmann et.al [52], Zhang et.al [35], Amin et.al [45]
Well-Being Assessment (WBA) Zhang et.al [35], Amin et.al [45]
Engagement Measurement (EM) Zhang et.al [35], Amin et.al [45]
AG Emotional Dialogue (ED) Empathetic Response Generation (ERG) Sibyl [53], PromptMind [54], PEGS [55], Chain of Empathy [56], ECoT [57]
Emotional Support Conversation (ESC) SoulChat [58], EmoLLM [59], PICA [60], Cue-CoT [61]
Review Summarization (RS) Opinion Summarization (OS) TCG [62], LASS [63]
2.1 Affective Understanding
2.1.1 Sentiment Analysis
AU tasks aim to enhance the model’s ability to recognize and understand various emotions, with Sentiment Analysis (SA) being the core task. The most fundamental task within SA is the Polarity Classification (PC) tasks [64], which generally requires determining whether the sentiment of a given text is negative, positive, or neutral. However, since PC tasks only divide emotions into two or three categories, more detailed Emotional Classification (EC) tasks [65, 66] with more labels have emerged to distinguish emotions better.
Another critical task is Aspect-Based Sentiment Analysis (ABSA) [67, 68], which differs from PC and EC by focusing on sentiment analysis for more specific aspects or attributes. To achieve more fine-grained sentiment analysis, ABSA involves a series of subtasks. The foundation lies in identifying specific aspects or attributes from the text, a process known as Aspect Term Extraction (ATE) [69, 70]. Once these specific aspects are identified, the next step is to perform sentiment analysis for each aspect, known as Aspect-Oriented Sentiment Classification (ASC) [71, 72]. This task aims to determine the user’s sentiment towards each aspect. To further enhance the efficiency and accuracy of the analysis, researchers have proposed Joint Aspect-Sentiment Analysis (JASA) [73]. This method combines aspect extraction and sentiment classification through joint modelling, better capturing the relationship between aspects and sentiments, and enhancing overall performance. Building on this, the Aspect-Category-Opinion-Sentiment (ACOS) task [74] further expands the dimensions of analysis. ACOS considers not only aspects and sentiments but also categories and specific opinions.
In addition to more fine-grained research on emotion labels and aspects, SA tasks also explore the quantification of emotion intensity, known as Emotional Intensity Detection (EID) [75, 76]. EID not only identifies the type of emotion but also assesses its intensity. Furthermore, considering that human emotions are not always explicitly expressed in the text, researchers have proposed the task of Implicit Sentiment Analysis (ISA) [77, 78]. ISA aims to identify and analyze the implicit sentiments in text that are not conveyed through explicit emotional vocabulary or expressions. Moreover, recognizing that emotions can be expressed through various modalities beyond text, researchers have introduced Multimodal Sentiment Analysis (MSA) tasks [79, 80], which integrate textual, visual, and acoustic information to provide a more comprehensive understanding of sentiment.
From the initial SA to more fine-grained ABSA, EID, and ISA, and further to MSA, SA tasks continuously evolve towards more profound, comprehensive, and accurate directions to better understant and apply human emotions.
With the continuous development of human-computer interaction technologies [81], Emotion Recognition in Conversation (ERC) [65, 82, 83] has become increasingly critical. The more challenging task, Emotion and Intent Joint Understanding in Multimodal Conversation (MC-EIU) is proposed. MC-EIU aims to decode the semantic information manifested in a multimodal conversational history while simultaneously inferring the emotions and intents for the current utterance [84]. The task of ERC requires identifying the emotions expressed in each utterance and tracking the emotional dynamics within the context of the conversation to achieve a comprehensive understanding of these dynamics. In a related but distinct direction, the Emotion Cause Pair Extraction (ECPE) [85] focuses on extracting emotions and their corresponding causes from the text. This task necessitates accurately identifying emotional expressions and pinpointing the specific causes of these emotions.
2.1.2 Subjective Text Analysis
While significant progress has been made in SA and ECR within AC, researchers are simultaneously exploring the potential of Subjective Text Analysis (STA) to provide even deeper insights into user experience and perspective nuances. This burgeoning field of study enriches by delving into the complexities of user intention and attitude revealed through textual information.
In particular, STA encompasses a variety of tasks, such as Personality Assessment [86], Suicide Tendency Detection [87], Toxicity Detection [^1] https://kaggle.com/competitions/jigsaw-toxic-comment-classification-challenge, Sarcasm Detection [88], Well-being Assessment [89], and Engagement Measurement [^2] https://www.kaggle.com/datasets/thedevastator/social-media-interactions-on-tedtalks-dataset, all of which aim to analyze subjective emotions, opinions, and attitudes within text. Compared to classic SA, these tasks receive less attention and fewer research resources.
2.2 Affective Generation
2.2.1 Emotional Dialogue
Affective Generation (AG) aims to enhance a system’s capacity to respond to and engage with various emotions. Within this field, Empathetic Response Generation (ERG) [11] and Emotional Support Conversation (ESC) [12] represent two prominent research areas. ERG tasks focus on training models to comprehend a conversation’s content and emotional nuances, enabling them to generate appropriate and empathetic responses. The emergence of multimodal data and research has further propelled this field, with researchers recognizing the efficacy of multimodal responses in fostering empathy within these models, leading to the development of Multimodal Empathetic Response Generation (MERG) [80, 55]. In contrast, ESC tasks adopt a more targeted approach to provide users with professional emotional support and assistance in navigating emotional challenges. Consequently, evaluating emotional support conversation tasks necessitates considering traditional dialogue evaluation metrics and the effectiveness of the emotional support provided [12, 66].
2.3 Review Summarization
Review Summarization (RS) is a broader task, and opinion summarization is a core component. Opinion Summarization (OS) [90, 91] is the automated process of condensing multiple opinions on a specific topic, typically from social media and user-generated content, into concise and meaningful summaries. The goal is to help users quickly understand a topic’s overall sentiment and key points without reading through large volumes of text. However, it faces challenges due to the unstructured and noisy nature of social media content and issues like sarcasm and diverse language use.
It is worth noting that research on AG lagged behind AU in the pre-LLM era due to limitations in model capabilities and technology. However, with their exceptional sequence generation abilities, the advent of LLMs has ushered in new possibilities for affect generation, making it a focal point of research in the LLM era [92].
3 Preliminary Study
Large Language Models (LLMs), with their powerful natural language processing, are rapidly transforming the field of Affective Computing. This section explores the emerging applications of LLMs in crucial areas of Affective Computing: Affective Understanding and Affective Generation. We review the current state of research, highlighting the strengths and limitations of LLMs in these tasks and suggest promising directions for future exploration.
3.1 Affective Understanding
3.1.1 Sentiment Analysis
Sentiment Analysis (SA), a classic task in the Affective Computing domain, focuses on the automatic identification and categorization of emotions and sentiments expressed through contents from users [93, 94, 79].
To evaluate the performance of LLMs in SA tasks, researchers conduct assessments on multiple datasets and discover that the performance of LLMs is closely related to the complexity of the SA tasks they face. For more straightforward sentiment analysis tasks, such as Polarity Classification (PC), closed-source large language models like GPT-3.5-turbo demonstrate robust zero-shot learning capabilities and can deliver satisfactory results [24, 95, 96, 97, 98]. However, when confronted with complex SA tasks requiring deep semantic understanding or structured emotional information, such as Emotional Classification (EC), Aspected-Based Sentiment Analysis (ABSA), Emotional Recognition in Conversation (ERC), etc., which involves identifying multiple sentiments corresponding to different aspects of a post, or Emotion Cause Analysis, the performance of LLMs still lags behind that of pre-trained language models (PLMs) fine-tuned on specific datasets [95, 99, 100, 101]. This discrepancy may stem from LLMs lacking training data tailored to specific tasks and having insufficient capabilities to process complex semantic information.
Although the zero-shot learning capabilities of open-source LLMs are inferior to those of closed-source models, their superior few-shot learning abilities give them an advantage on datasets with limited training data or imbalanced distributions, outperforming PLMs in such scenarios. However, when data is abundant and label distribution is balanced, PLMs can still perform better due to their fine-tuning advantage on specific datasets [102]. Notably, LLMs show superior generalization capabilities in SA tasks compared to PLMs. By simply altering the prompts, LLMs can adapt to different datasets and tasks without requiring model retraining [100].
However, the performance of LLMs is also affected by prompts and decoding parameters. For example, minor text attacks like synonym substitution can lead to fluctuations in their performance. Additionally, because of the auto-regressive decoding mechanism, parameters such as temperature and Top-P can significantly impact the effectiveness of LLMs in sentiment analysis tasks. Typically, a lower temperature value (below 0.3) and a lower Top-P value are necessary for achieving better results [19, 103]. This suggests that although research shows carefully crafted prompts can greatly improve LLMs’performance, developing optimal prompts and setting effective hyper-parameters for specific tasks and contexts remains an unresolved research challenge.
3.1.2 Subjective Text Analysis
LLMs demonstrate considerable performance variations across various Subjective Text Analysis (STA) related tasks. For instance, they excel in tasks involving negative emotion recognition, such as well-being assessment and toxicity detection. The superior performance could be due to the focus on safety and value alignment during the Reinforcement Learning from Human Feedback (RLHF) stage of LLM training [104, 19, 99]. However, in tasks requiring the understanding of implicit emotional cues, such as engagement measurement, personality assessment, and sarcasm detection, LLMs exhibit a performance gap compared to PLMs fine-tuned on task-specific datasets. This discrepancy may stem from a need for more specialized training data during the supervised fine-tuning (SFT) phase of LLM development. Optimizing prompts for LLMs can enhance their performance on these specific tasks [19, 99]. Furthermore, leveraging their robust capacity for capturing contextual information, LLMs consistently demonstrate superior performance in processing longer texts compared to shorter texts across various subjective text analysis tasks [99, 24]. LLMs perform slightly worse than SFT models tailored to specific tasks in complex sentiment computing areas like sarcasm detection and subtle sentiment analysis. However, the substantial computational and time resources needed to fine-tune these large language models present a significant hurdle. Consequently, developing LLMs with outperformance and minimal training for specific tasks is an important research priority.
3.2 Affective Generation
3.2.1 Emotional Dialogue
Empathetic Response Generation (ERG) and Emotional Support Conversation (ESC) are two prominent tasks in AG that attract significant attention, utilizing the EMPATHETICDIALOGUES [11] and ESConv [12], respectively. Early research suggests that LLMs possess immense potential in emotionally-aware generation tasks. For example, GPT-3.5-turbo has shown competitive performance across multiple datasets [4], while GPT-4 generates empathetic responses in various emotional scenarios. Human evaluations indicate that GPT-4’s responses exceed human performance by an average of 10% in empathy levels [105].
Unlike earlier models, LLMs produce longer and more diverse text [105], introducing new evaluation challenges. Traditional automatic evaluation methods based on overlap metrics need to adequately measure the performance of LLMs in emotional generation tasks, while manual evaluation is costly and burdensome to scale. Developing novel evaluation methods and establishing standardized frameworks to accurately measure LLM-generated affective responses’ quality, empathy, and emotional intelligence is crucial for advancing research in this area.
3.2.2 Emotional Intelligence
Emotional Intelligence (EI) [106, 107, 108] is the ability to recognize, understand, manage, and use emotions effectively in various contexts. In the era of LLMs, assessing and enhancing EI in LLMs is an emerging challenge.
EI is crucial for enhancing the naturalness and effectiveness of human-computer interaction and is a vital component in achieving Artificial General Intelligence (AGI). To explore whether LLMs possess EI, researchers undertake investigations from various perspectives. Current studies[109, 110] draw upon assessment methods from the field of human psychology, employing tools, like the Levels of Emotional Awareness Scale (LEAS) [111], to evaluate the ability of LLMs to perceive emotions. LLMs can generate relatively appropriate emotional responses, which has been evidenced.
However, using human-centric evaluation systems to measure the EI of LLMs has limitations, including the fact that the scales are designed with humans in mind, not explicitly specifically for LLMs and that the evaluations rely on human experts to conduct the assessments. Researchers have developed evaluation tools designed explicitly for LLMs to address these issue, such as the Situational Evaluation of Complex Emotional Understanding (SECEU)[112]. Moreover, to improve the objectivity and efficiency of evaluations, EQ-bench has been proposed [113]. It removes the need for expert assessments and instead directly evaluates the ability of LLMs to understand complex emotions and social interactions by predicting the intensity of emotional states in dialogue characters. While LLMs have demonstrated remarkable performance on SECEU and EQBench, which focus more on the cognition and understanding of human emotional scenarios, there is still insufficient evidence to suggest that they possess emotional intelligence comparable to humans. To address this gap, researchers propose a new task called Sentiment and Opinion Understanding of Language (SOUL), which aims to comprehensively evaluate the model’s ability to understand complex emotions through two sub-tasks, including Review Comprehension and Justification Generation. However, experimental results reveal a substantial performance gap of up to 27% between LLMs and human performance on the SOUL task[114]. It highlights the considerable room for improvement in the capacity of LLMs to understand the nuances and complexities of human emotions.
4 Instruction tuning
LLMs show excellent performance on different tasks, such as Question Answering, Sequence Generation, and so on [115]. Instruction Tuning [116] is introduced to adapt LLMs to specific tasks, thereby maintaining the benefits of the original model while improving the performance of LLMs on specialized downstream tasks.
Due to the substantial size of LLMs, which often exceed ten billion parameters, Full Parameter Fine-Tuning (FPFT) of all parameters of the complete model could be more practical. To reduce the cost of model training, a series of Parameter-Efficient Fine-Tuning (PEFT) approaches are proposed, such as LoRA [117], Prefix Tuning [118], and P-Tuning [119], as shown in Figure 2. LoRA reduces the number of parameters that need to be trained during the fine-tuning process by freezing all the original parameters and injecting a pair of low-rank decomposition matrices next to the original weights. Prefix Tuning adds trainable "prefix" tokens related to specific tasks as an implicit learnable prompt to the input sequence of the LLM. It involves inserting trainable free vectors into each layer of the Transformer. Applying PEFT in instruction tuning can effectively reduce the training costs in AC tasks.
(a) Full-Model Tuning
(b) Multi-LoRA
(c) LoRA
(d) Prompt Tuning
Figure 2: Illustrations of the instruction tuning methods. Pink blocks denote trainable modules, blue blocks signify frozen modules, and orange blocks represent optionally added modules. Subfigure 2(a) depicts the FPFT method, the transformer blocks of two distinct model architectures are fine-tuned. Subfigure 2(b) introduces the multi-LoRA method. In this approach, only the added multi-LoRA models are trained. This method allows for the optional incorporation of router modules to manage the weights of each LoRA module. Subfigure 2(c) demonstrates the LoRA method, The traditional LoRA method involves the addition of only two low-rank matrix modules, LoRA_A and LoRA_B. VeRA [120] allows the addition of two trainable vectors, d and b, augmenting the LoRA layer. This approach also shares LoRA_A and LoRA_B’s parameters to every LoRA layer. Moreover, when the rank of LoRA in each layer varies, this method can transition into adaLoRA [121]. Subfigure 2(d) illustrates the soft prompt tuning, where a trainable prefix is appended to the model’s embedding layer, known as p-tuning. Furthermore, scientific prefixes can be optionally added to each layer of the model, with only the added prefixes being trained, a technique referred to as prefix-tuning.
4.1 Affective Understanding
In addressing affective computing tasks, PLMs typically necessitate separate approaches for affective understanding and generation tasks. In contrast, LLMs can unify all affective computing tasks into a sequence generation format. However, directly employing LLMs for AU tasks does not offer significant advantages over fine-tuned PLMs. Consequently, numerous recent studies focuse on constructing task-specific instruction tuning templates tailored to the unique characteristics of various tasks. These studies apply instruction tuning to the models, aiming to enhance the performance of LLMs on specific tasks or within particular domains.
4.1.1 Polarity & Emotional Classification
Polarity Classification (PC) and Emotional Classification (EC) tasks are frequently employed to assess a model’s comprehension and perceptual abilities. Several studies pioneer the exploration of the effects of instruction tuning of LLMs on the SA task [122, 123]. The T5[124] series model, in particular, demonstrates an improved understanding of the sentiment content of text after instruction tuning. Recent research [123] focuses on exploring the different PEFT methods, adopting the LoRA and P-tuning methods to fine-tuning ChatGLM2[125]. Experimental results demonstrate that after instruction tuning, LLMs are greatly improved in eight datasets compared with the model without instruction tuning. However, the experimental results must conclusively determine which instruction tuning method is more effective in enhancing performance on PC. The emotional polarity of financial data is also a PC task. A recent study proposes a retrieval-augmented LLMs framework [126] consisting of instruction-tuned LLMs and a retrieval-augmented component. This method can classify the sentiment of data in the financial domain.
While these studies demonstrate LLMs’ effectiveness in PC and EC tasks, it is equally important to understand their underlying sentiment processing mechanisms. Research shown that words with sentiment information can influence the overall sentiment of the text [38]. Building on this understanding, LLMs leverage techniques of instruction tuning to grasp word polarity, enabling them to effectively perform methods PC and EC tasks.
Although these studies primarily focus on text-based PC and EC tasks, it is essential to recognize that real-world user expressions are not limited to text alone. User-generated content often encompasses multiple modalities, including images and audio, which can provide additional context and nuance to sentiment analysis. WisdoM [39] leverages contextual world knowledge derived from LMM to enhance multimodal sentiment analysis (MSA). Additionally, the authors introduce a training-free module called "contextual fusion," which can reduce noise in the contextual data and improve model classification accuracy. WisdoM significantly outperforms existing state-of-the-art methods on the MSED dataset, demonstrating its effectiveness in integrating contextual knowledge to improve emotion classification. Despite these promising advancements, it is worth noting that research on MSA tasks based on LLMs or Multimodal Large Language Models (MLLMs) still needs to be improved.
4.1.2 Aspect-Based Sentiment Analysis
Aspect-Based Sentiment Analysis (ABSA) encompasses multiple subtasks, such as Aspect Term Extraction (ATE), Aspect-oriented Sentiment Classification (ASC), Joint Aspect-Sentiment Analysis (JASA), and Aspect-Category-Opinion-Sentiment (ACOS), and others. The difficulty of these subtasks progressively increases, starting with known aspect items, then identifying aspect items, and finally performing the corresponding sentiment analysis. Traditional PLMs often design a separate module for each ABSA subtask, which reduces the correlation between ABSA subtasks. LLMs can leverage their strong generalization abilities to construct a unified framework for all ABSA subtasks.
[…]
4.1.3 Emotion Recogniztion in Conversation
Emotion Recogniztion in Conversation (ERC) tasks involve recognizing the sentiment label of each utterance in the conversation, which is highly dependent on the conversation history and the speaker.
InstructERC addresses these challenges by developing a template that includes a system prompt, dialogue history and the label map. It also introduces Instruction Tuning using LoRA for ERC tasks [49]. Additionally, InstructERC creates two auxiliary tasks, speaker identification and emotion impact prediction, to enhance the model’s understanding of speaker information. Building on this approach, CKERC [48] also employs LoRA to solve ERC tasks. It further leverages the LLaMA2-chat model [130] to generate common sense knowledge, enabling a more comprehensive understanding of the model of dialogue context. This model achieves state-of-the-art results on the MELD [131] and EmoryNLP [132] datasets.
Researchers extend ERC tasks to incorporate multimodal inputs, considering the frequently multimodal aspects of human communication, leading to the development of Multimodal Emotion Recognition in Conversation (MERC). DialogueLLM [133] generates text descriptions of visual information through GPT-4, combines these descriptions with the conversation history as an instruction dataset, and uses LoRA to train LLaMA2. Through experiments, it has been found that adding information from other modalities can improve the model’s ability to understand dialogue information.
Due to the high similarity of mission objectives and approach between the SA and ERC tasks, researchers try to build a unified model that can accomplish both tasks simultaneously. Emollms [134] develops a multi-task instruction dataset, AAID, covering SA and ERC. Experimental results show improved performance on each task after multi-task training. Similarly, Unimse [47] introduces a multimodal affective knowledge-sharing framework, unifying MSA and MERC tasks and fusing information from multiple modalities through the T5 model. Notably, it is the first method to solve MSA tasks in a generative method. These studies indicate that mixing datasets from multiple tasks can improve a model’s generalization ability, enabling performance on untrained datasets comparable to fine-tuned PLMs.
4.1.4 Emotion Cause Extraction
Emotion Cause Extraction (ECE) has garnered widespread attention in recent years. This task aims to extract potential causes behind certain emotions in text, but its application scenarios are significantly limited due to the need for annotating emotional information. To address this issue, the Emotion-Cause Pair Extraction (ECPE) task propose. The ECPE task simultaneously extracts potential Emotion-Cause pairs and their corresponding causes within a document, thereby reducing the dependence on emotion labels [135].
Most current research on ECPE tasks involves fine-tuning traditional PLMs [136, 137] or leveraging the understanding capabilities of LLMs using well-designed prompts [50]. These methods primarily focus on the Textual Emotion-Cause Pair Extraction in Conversations task (TECPE). However, emotion causes may also be embedded in other modalities, leading to the development of Multimodal Emotion-Cause Pair Extraction in Conversations task (MECPE) [138]. LLMs shows potential in better integrating information from multiple modalities for such tasks.
To address the MECPE task, NUS-Emo [139] proposes a dual-component system that operates in two stages. In the first stage, the system uses ERC data to fine-tune the LLMs, enhancing its emotion recognition capability. It then employs ImageBind to encode image information fused with the recognized emotional information. In the second stage, the system inputs the emotional information into the model for inference and uses the inferred dataset as the training set for further fine-tuning. This iterative fine-tuning approach aims to improve the model’s ability to infer emotion causes across multiple modalities.
4.1.5 Subjective Text Analysis
LLMs are employed to explore the Stance Classification task. In this task, the model determines whether a user agrees, disagrees, or is neutral on a given point of view [140]. The study investigates ten open-source models and seven prompting schemes, revealing exciting conclusions. These include the fact that LLMs are competitive with in-domain supervised models and that the fine-tuning process only sometimes leads to better performance than zero-shot models. This is likely because the fine-tuning makes the model too specialized, thus unable to generalize to out-of-domain data points and varying stance definitions.
4.2 Affective Generation
In the current landscape of artificial intelligence, content generation that expresses sentiment is applied to various practical fields, including emotional dialogue, chatbots, emotional support conversation, and more [5]. It enhances humanized services by analyzing and generating emotional information from users.
4.2.1 Emotional Dialogue
Emotional Dialogue (ED) refers to the ability to understand and share another person’s feelings and then respond in a way that conveys this understanding and compassion. LLMs must to possess high emotional intelligence to resonate with users in their replies. However, due to the datasets’ limited and relatively homogeneous nature, the generated responses often tend to be formulaic and mechanical, resulting in lower-quality interactions. Researchers aim to construct more diverse empathetic dialogue datasets. Considering the high labor costs of manually constructing datasets, current research primarily employs synthesis data for LLMs methods to build high-quality empathetic dialogue datasets [141, 53, 55].
Several studies use these methods to create diverse datasets. For example, the ExTES dataset is expanded using GPT-3.5-turbo and self-instruct [142], resulting in a total of 11,177 conversations where users display negative emotions and the assistant provides more positive support responses [141]. Soulchat [58] employs GPT-3.5-turbo to expand existing dialogue datasets through iterative rounds, using prompts to generate multi-turn conversations. Another notable example is STICKERCONV, a multi-scene, multimodal empathetic dialogue dataset built by generating user information from a multi-agent system [55]. Data generation through LLMs can significantly reduce labor costs. The dialogue scenarios and content are wider than providing more diverse data types. However, the quality of the generated data depends on the model’s capabilities and may need to be more consistent, often requiring secondary screening. To mitigate quality issues, some studies avoid the direct generation of dialogue data and instead generate objective knowledge, thus avoiding the quality problem of the generated data. A recent study [53] proposes a novel approach using GPT-3.5-turbo to generate common sense knowledge about conversations. This generated knowledge encompasses the discourse’s underlying reasons, emotions, and intentions. By incorporating this knowledge into the model’s training process, the research enhances the model’s ability to comprehend the deeper meaning of conversations and generate more empathetic responses. Experimental results demonstrate that each type of knowledge is adequate for empathetic response generation.
4.2.2 Chatbot
Current datasets still limit most of the ERG and ESC tasks described above and need to capture the diversity of users and scenarios in real life fully. Chatbots capable of understanding user input and emotions are designed to address the practical application of emotional dialogue systems. These chatbots can offer encouragement during adversity and share in users’ happiness.
PICA [60] and EMOLLM [59] are two notable examples of such chatbots. While LLMs can function well as "assistants," their emotional intelligence is often lacking, and long replies frequently fail to convey empathy. To address this issue, PICA is proposed. It uses P-tuning to fine-tune the ChatGLM2-6B model on two synthetic Chinese empathetic dialogue datasets, effectively enhancing the emotional capabilities of LLMs. EMOLLM integrates a series of LLMs to support and assist users with mental health counseling tasks. It uses the QLoRA[143] method to fine-tune multiple LLMs, such as Qwen-7B, LLaMA3-8B-Instruct, and others, across different languages. By supporting various roles, EMOLLM enables the chatbot to generate more contextually appropriate user responses.
4.3 Multi-task of Affective Computing
Many experiments show that LLMs possess strong comprehension and generalization capabilities. Furthermore, their sequence generation ability enables them to construct a unified paradigm for AU and AG tasks. Consequently, researchers began exploring the development of a general AC model.
However, current fine-tuning methods present limitations when dataset sources differ across tasks. Fine-tuning a model under these conditions can create conflicts and impact the original world knowledge embedded in the model [144]. These challenges highlight the need for more research on ensuring high-quality performance across all sentiment analysis tasks.
To address these issues, researchers propose a novel Modular Emotional Intelligence enhancement method (MoEI) featuring two collaborative techniques—Modular Parameter Expansion (MPE) and Intra-Inter Modulation (I2M)—is proposed to fit most emotion analysis tasks [145]. In conjunction with this method, the study develops EIBENCH, a unified affective computing dataset containing 88 datasets across 15 EI-related tasks, is introduced. MoEI adopts the Mixture-of-LoRA (MoLoRA) method for fine-tuning to improve EI while maintaining the generative ability of the model.
In multimodality, MEILLM[146] integrates deep feature vectors extracted from visual, textual, and audio models into baichuan13B-chat through the MLP layer. After fine-tuning the instructions, he can complete a variety of emotional interaction tasks, including psychological assessment, psychological portrait, psychological counseling, etc.
4.4 Challenge
While Instruction Tuning has enabled LLMs to make significant strides in AC tasks, several challenges remain to be addressed.
  • FPFT LLMs for specific domains are time-consuming, labor-intensive, and requires significant computational resources. While PEFT can reduce resource consumption in fine-tuning LLMs, they are more suitable for specific datasets and often sacrifice the generalization ability of LLMs. There is a need for scholarly research about the comparative efficacy of various PEFT methods on AC tasks. Consequently, developing a simple yet effective fine-tuning approach that balances efficiency, task-specific performance, and generalization capability remains a significant challenge in LLMs.
  • While LLMs with instruction tuning demonstrate excellent performance in AC tasks, they are still slightly inferior to the previous fully fine-tuned PLMs method in some tasks. Achieving state-of-the-art performance with minimal training remains a major challenge.
  • Despite their known universality, LLMs often struggle to simultaneously excel in both AU and AG within the domain of AC. Developing a unified model that effectively addresses AU and AG remains a significant challenge.
  • Research in niche areas of AC tasks, such as financial sentiment analysis, sarcasm detection, and metaphor analysis, still needs to be improved. Furthermore, the full potential of LLMs in these specialized AC tasks has yet to be thoroughly explored and evaluated.
5 Prompt Engineering
LLMs, owning powerful understanding and generation capabilities, pave new ways for tasks associated with AC. Despite pre-training or supervised fine-tuning (SFT), effectively utilizing LLMs for AC remains challenging. Prompt engineering [27] achieves higher accuracy and reliability in the AC tasks by designing appropriate input prompts that guide LLMs to generate the desired outputs. This section introduces the latest advancements in prompt engineering for AC, focusing on key technologies and application cases to provide references and insights for related research. Current research identifies four main prompting techniques, as shown in Figure 3: Zero-shot Prompt, Few-shot Prompt, Chain-of-Thought (CoT) Prompt, and Agent.
Zero-shot Prompt [27] is a method where the model generates text about a specific topic or domain without any demonstrations or specific instances from training data. This approach is structurally simple, easy to understand, and capable of quickly adapting to different fields and tasks. By leveraging model’s extensive knowledge and intrinsic reasoning abilities, zero-shot prompts can effectively generate the required text. This makes them particularly effective in applications where high-quality content needs to be generated quickly, especially when data is insufficient or difficult to obtain.
Few-shot Prompt [27] provides models with a few examples or demonstrations related to the given task. These examples and the task description, are input as prompts to help the model better understand the specific task. The model can perform efficient reasoning and generate high-quality output by learning from a few examples and following the task instructions. Few-shot prompt improves the performance of many complex tasks and is particularly advantageous when there are limited data samples or data imbalances.
Figure 3: The Overview of Prompt Engineering
CoT Prompt [147] is a technique that significantly enhances the reasoning ability of LLMs. Instead of directly prompting the LLMs to generate output, the CoT prompting method uses a step-by-step reasoning process to guide the models through logical thinking. This method involves explicitly outputting intermediate steps as part of the reasoning chain. CoT prompting allows LLMs to generate more accurate and contextually relevant responses. It is effective in various complex reasoning tasks, such as arithmetic, common sense, and symbolic reasoning. Additionally, the intermediate steps make it easier for users to understand the thinking process of a model, thus improving the interpretability of the reasoning of LLMs.
Agent [92] is an emerging and promising prompt engineering technology. It functions as an intelligent agent capable of perceiving its environment, understanding it autonomously, making decisions, and executing actions. It can independently think and call upon various tools to gradually achieve goals. In an LLM Agent, the LLM acts as the agent’s brain, collaborating with multiple LLM components to complete each part of a task. The profile, tool, module, plan, and action modules perform their respective tasks based on the prompt method. The LLM brain then synthesizes the results from each module to achieve better outcomes. This collaborative approach makes agents particularly effective in solving complex tasks that require multiple steps or diverse knowledge. They are often employed in scenarios that demand role-playing [92] or coordinated actions [148], showcasing their versatility in handling intricate, multi-faceted challenges [149].
While these prompting techniques share some commonalities, they have distinct characteristics and applications. The Zero-shot Prompt method guides the model using a sentence or paragraph as a prompt text. The other three methods build upon this with various modifications and expansions. The Few-shot Prompt method adds task-solving examples to the prompt texts to assist in task completion. The CoT Prompt method employs multi-step chain thinking, using different prompt texts to complete the task at each step. CoT can also be combined with the few-shot method, using CoT’s the multi-step thinking process of CoT as examples to guide the generation of the thought chain. The agent method differs from the previous three methods in adopting a collaborative approach with multiple LLMs. Each LLM can use any prompt form, such as zero-shot, few-shot, or CoT, and the information output by each LLM is combined to complete the task. Figure 3 illustrates the similarities and differences among the four methods.
5.1 Affective Understanding
5.1.1 Zero-shot for Affective Understanding
In the field of affective understanding, zero-shot prompting is initially applied. Researchers use descriptions of sentiment analysis tasks as prompts, placing them before the target text to guide LLMs in generating sentiment for the target text. Through this method, researchers can observe the resulting sentiment analysis outcomes, thereby gaining a more comprehensive understanding of the performance and potential of various LLMs in AU [45, 150, 151, 42].
To enhance the performance of LLMs in AU, researchers develop several advanced prompting techniques that go beyond simple task descriptions. One such approach is a role-playing zero-shot prompting method [40], where LLMs are assigned a specific role, such as an emotion expert, to perform sentiment analysis tasks. By assigning roles, the emotional capabilities of the model can be better utilized. Additionally, a straightforward yet effective approach is called EmotionPrompt, a series of emotion stimulus questions based on psychology are incorporated into the original task prompts [63] to explore the sentiment analysis capabilities of LLMs by embedding specific information in the prompts. In Speech Emotion Recognition (SER), researchers incorporates information from acoustic feature descriptors into the prompts [152], combining numerical acoustic features into the prompts to identify the underlying emotions conveyed in the speech.
Although these zero-shot prompting methods using fixed prompt content achieve notable performance, their flexibility and diversity are limited due to the constraints of the prompt content. Therefore, in the field of multimodal sentiment analysis, researchers explore the use of LLMs to generate prompts based on specific needs [39], and then use these generated prompts as inputs to guide LLMs in analyzing the sentiments involved in image-text pairs.
Additionally, more attention must be paid to interpretation or emotional response in a range of subjective tasks such as Metaphor Recognition [153], and Dark Humor Detection [52]. RiC (Reasoning in Conversation) [154] is a Zero-shot approach that addresses these subjective tasks through dialogue simulation. RiC extracts useful contextual information by simulating conversations, providing the underlying knowledge behind the dialogue for the final answer.
5.1.2 Few-shot for Affective Understanding
Few-shot prompting is introduced to handle more challenging tasks, such as complex Aspect-Based Sentiment Analysis (ABSA) and Multimodal Sentiment Analysis (MSA).
For the challenging Aspect-Sentiment Quad Prediction (ASQP) [155] task, the Self-Consistent Reasoning-based Aspect-Sentiment Quad Prediction model, named SCRAP, is proposed [44]. This model uses prompt examples to generate reasonable reasoning-related content, known as few-shot CoT prompting, for subsequent fine-tuning. By leveraging a few reasoning strategies as prompts to enhance LLM reasoning and integrating the generated reasoning into the subtasks of ABSA, the ability of the model to handle complex reasoning tasks is improved, along with its interpretability and accuracy.
Few-shot Prompt method is also applied to MSA in AU. WisdoM is proposed to address four Multimodal Entity-Based Sentiment Analysis (MEBSA) subtasks [156]. WisdoM converts visual inputs into textual counterparts regarded as auxiliary sentences and selects positive and negative instances based on text similarity to construct examples. In the context generation stage, LVLMs [157, 158] use the prompt templates generated in the prompt templates generation stage to generate context that explicitly contains world knowledge based on a given image and sentence demonstration. Although WisdoM is a straightforward and effective plugin framework that significantly enhances the performance of multimodal sentiment analysis, the rich information in visual content still need to be fully utilized. Therefore, compared to the easy-to-understand semantic information from the text modality, fully extracting and utilizing the content from the image modality with multimodal LLMs remains a significant challenge.
Additionally, in the Depression Detection task, Few-shot Prompt method guides data generation to enhance the Structural Element Graph (SEGA) [159] further. Researchers use explicit principles to guide the synthesis process and provided LLM with data generation examples to obtain high-quality synthetic data.
5.1.3 CoT for Affective Understanding
The information provided by zero-shot or few-shot prompting through a single input-output process must be more sufficient to support the completion of complex AU tasks. Therefore, researchers explore the CoT method, which incorporates a series of intermediate reasoning steps during the prompting process to enhance the model’s understanding of inputs, thereby better grasping the emotions and nuanced meanings.
Due to the complexity of affective understanding, introducing information that aids in affective understanding can significantly enhance the model’s ability to perceive and process emotions. Therefore, researchers propose a three-hop reasoning chain framework called THOR [46] for Implicit Sentiment Analysis (ISA). Based on the importance of hidden opinions in conversations, the THOR CoT framework sequentially extracts fine-grained aspects, analyzes the latent opinions on these aspects, and derives accurate implicit sentiments based on this information.
Unlike conventional methods that directly extract regular information from known texts, the Chain of Empathy (CoE) [56] adopts a psychological perspective, incorporating various psychotherapy methods such as Cognitive Behavioral Therapy (CBT) [160], Dialectical Behavior Therapy (DBT) [161], Person-Centered Therapy (PCT) [162], and Reality Therapy (RT) [163]. It designs four thinking chain methods to interpret user’s psychological states of users from a professional standpoint. Ultimately, CoE uses the information from these psychotherapy methods to infer human emotional states,achieving a more accurate understanding and processing of emotion. Drawing inspiration from the Stimulus-Appraisal-Emotion process within the cognitive appraisal theory of emotion generation, the ECR-Chain framework [51] undertakes a multi-step approach. First, the ECR-Chain summarizes the dialogue topics. Next, it generates speaker responses to these specific topics using LLM. Subsequently, the model infers the user’s target appraisal and extracts stimulus information from the dialogue. Finally, based on this information, the ECR-Chain infers the underlying cause of the emotion.
Methods such as COE and ECR-Chain generate beneficial inference information based on task requirements and achieve good results. However, they face significant challenges in complex extraction tasks, such as Joint Aspect Sentiment Analysis (JASA) and Emotion-Cause Pair Extraction (ECPE). A multi-step reasoning approach is proposed to decompose the tasks. A multi-step reasoning method is proposed to decompose the task. Specifically, to improve the performance of sentiment analysis (SA), RP-CoT [40] first identifies the sentiment polarity information corresponding to various aspects of the text instance and then classifies the overall sentiment polarity of the entire text based on the obtained fine-grained sentiment information. Similarly, when performing the task of ECPE, DECC [50] first identifies the emotions present in the given context and then locates and reserves the specific utterances (emotional clauses) associated with each emotion. Finally, for each emotion, the most probable clause is selected as its cause, and the pairs of emotional and cause clauses are outputted.
In addition to using CoT to generate information that aids in understanding emotions and decomposing complex tasks, the CoT method is also applied to prompt construction. The Evolutionary Multi-Objective (EMO) process [164] inputs manually crafted original prompts into LLMs and evaluates their fitness through generated stories. EMO then randomly selects two prompts from the original set and recombines them using crossover operators. Applying a new set of mutation operators to the recombined prompts, it ultimately produces new prompts that are better suited to the task.
The CoT Prompt method can also be combined with Few-shot Prompt method. In the task of Market Sentiment Analysis [165], it is necessary to master financial market and social media terminology, making manual evaluation extremely challenging. The lack of high-quality labeled data limits the application of traditional supervised learning methods. Researchers have drawn on the CoT concept by constructing contextual examples and using Few-shot Prompt method to prompt LLMs to generate CoT summaries [166]. This guides LLMs to recall relevant financial domain knowledge beforeconclusions, thereby injecting relevant knowledge into Market Sentiment Analysis.
5.1.4 Agent for Affective Understanding
In previous studies, the standard paradigm for affective understanding relies on a single LLM and makes decisions [46, 156]. However, this method has a critical drawback: the output generated by a single LLM may not provide the optimal decision. Therefore, the Agent technology is further proposed to enable the collaboration of multiple LLMs.
The Generator-Discriminator Role-switching Decision-Making framework is proposed to achieve multi-LLM negotiation for sentiment analysis [41]. This framework consists of a reasoning-infused generator, responsible for providing decisions and rationale, and an explanation-deriving discriminator, which evaluates the credibility of the generator. The generator and discriminator iteratively communicate until a consensus is reached. By leveraging the complementary capabilities of the two LLMs, sentiment interpretation and correction can be better achieved. The PANAS framework [167] is designed with multiple agents to perceive different information. New natural language experiences are input into Agent 1, which summarizes and extracts insights from memories to create norms. Agent 2 conducts context understanding, highlighting the differences between the current situation and the norms. Finally, Agent 3 evaluates emotional responses based on the provided context understanding.
In the field of Financial Sentiment Analysis (FSA), HAD [168] summarizes five common types of errors in FSA tasks and instantiates five specialized agents based on this prior knowledge: the mood agent, the rhetoric agent, the dependency agent, the aspect agent, and the reference agent. Each agent focuses on a specific type of error that LLMs are prone to in given FSA tasks. Ultimately, the FSA results are obtained through a shared discussion that considers the outputs of all agents.
5.2 Affective Generation
Prompt engineering has also been introduced in the field of Affective Generation to improve performance across various tasks and scenarios. The following introduces four applications of prompt engineering techniques in Affective Generation.
5.2.1 Zero-shot for Affective Generation
To adapt prompts to the task characteristics of Affective Generation, researchers propose the Perspective-Taking method [169], which enhances the capability of LLMs to generate empathetic dialogues by enabling perspective shifts. Within the HEF framework [54], small empathy models are first trained to generate the emotion cause and the main emotional categories in the dialogue, which are then provided as input to the LLM. This approach gives the model more accurate information, resulting in more precise empathetic responses generated under zero-shot prompts. Researchers design the TCG pipeline [62] based on GPT-3.5-turbo for the task of opinion summarization. Through sentence opinion extraction, grouping, and clustering, the pipeline ultimately produces a summary of the long-form’s opinions.
In the emotion-conditioned text generation task, researchers optimizes prompts by iteratively adding, replacing, and deleting the original prompts [170]. The optimized prompts better achieve the expected emotions, resulting in improved emotional text generation. Since the output of LLMs largely depends on prompts, users needing more background knowledge might experience a diminished interaction with chatbots due to the absence of contextual understanding. To address this issue, PromptMind [171] generates multiple contextually relevant prompts during the conversation for users to choose from, thereby prompting LLMs to produce answers that better meet user expectations and enhance human-chatbot interaction.
5.2.2 Few-shot for Affective Generation
In Sibyl framework [172], researchers randomly select a sample as a demonstration to prompt the LLM to generate four categories of visionary commonsense related to the dialogue as data for subsequent training of the dialogue generation model. Sibyl improves the empathy generation ability of the LLM [53]. In mixed-initiative dialogue generation tasks like PersuasionForGood and Emotional Support Conversations, generating responses by setting different strategies as examples is also a way to use few-shot prompting in generation tasks [173].
Researchers indicate that generated responses utilizing prompt engineering demonstrate high quality and adherence to semantic controls [173]. However, in emotional support tasks, the inherent preference of LLMs for specific strategies can hinder effective support despite improving the robustness of predicting appropriate strategies. To address this, researchers incorporate strategy descriptions and randomly select 2-Shot examples into prompts to enhance the LLM’s understanding of each strategy [174]. LLaMA2-7B [14] is leveraged for planning the subsequent action, followed by LLM generation of responses aligned with this strategy. Finally, the number of examples within the prompt is expanded. Notably, results suggest that while providing an appropriate number of examples can enhance preference bias, simply increasing the number of examples may only sometimes yield better results.
5.2.3 CoT for Affective Generation
In AG, information and knowledge are crucial for producing high-quality conversations. To obtain more valuable information, CoT is also introduced to assist affective generation. User states, serving as underlying linguistic cues, can reveal hidden needs, guiding the model to generate more personalized and engaging responses. The Cue-CoT framework [61] breaks down the reasoning process into several consecutive steps, progressively generating user state information such as personality, psychological state, and emotions based on the context, fully perceiving the linguistic cues within the context. The final step is to generate responses based on the previous reasoning output. The EDIT framework [175] also enhances dialogue response generation by detecting implicit user intentions. First, EDIT generates open-ended questions related to the context to capture potential user intentions. Then, it searches for answers in domain-specific knowledge bases to acquire additional knowledge. Finally, it enhances response generation by explicitly integrating this additional knowledge.
Since Affective Generation requires models to handle emotion-related issues in dialogues more proactively, planning strategic information becomes essential when utilizing LLMs to obtain information. The ProCoT [176] framework guides LLMs to generate descriptive information about intermediate steps in reasoning and planning, analyzing the following action towards the dialogue goal through dynamic reasoning and planning. Responses are generated based on this planning information, endowing the dialogue system with proactive dialogue capabilities. The ECoT framework [177] utilizes LLM to sequentially execute five steps: understanding context, recognizing others’ emotions, recognizing self-emotions, managing self-emotions, and influencing others’ emotions, aligning with human emotional intelligence to generate more emotionally rich responses. Additionally, to mimic the processes of recognizing, understanding, and regulating emotions, ESCoT [178] generates emotional stimuli and formulates strategies for emotion regulation to facilitate response generation.
Additionally, in the metaphor generation task, GROUNDS naturally serve as the CoT connecting TENOR and VEHICLE [179]. Therefore, researchers use the labels in annotated datasets as examples to guide the model in generating logical CoT content as GROUNDS. Ultimately, the model leverages these GROUNDS to generate coherent and comprehensive VEHICLES, thereby improving the quality of the generated metaphorical expressions.
5.2.4 Agent for Affective Generation
As a new technology based on LLMs, Agents are gradually being applied in AG. The Agent4SC system [55] generates user information, enhances memory capabilities, and formulates response strategies through the cooperation cooperation of agents, realistically simulating human behavior in using stickers, thereby enhancing multimodal empathetic dialogue. The COOPER dialogue framework [180] coordinates multiple specialized agents, each focusing on specific aspects of dialogue goals, to generate emotional support and persuasive dialogue responses.
Additionally, in the multimodal domain, researchers study multimodal Conversational Health Agents (CHAs) [181] and design five types of agents as components to interact with users to generate audio responses based on user emotions. These studies indicate that Agent has excellent potential in AG, effectively improving dialogue’snaturalness and empathetic capability.
5.3 Challenge
Despite the significant achievements of Prompt Engineering in the field of AC, the following challenges persist:
  • As an efficient method with low resource requirements, it greatly reduces the cost of models. However, its performance indicators across various tasks still need to catch up to those of fine-tuned models, unable to meet practical application needs.
  • Currently, the most widely used and advanced CoT method has significantly improved the effectiveness of AC tasks. However, its underlying logic aims to fit human reasoning processes, which conflicts with the unique mechanisms of human emotion generation. This often results in noise during the intermediate reasoning process, affecting the final inference outcomes.
  • Moreover, since emotions are influenced by various factors such as psychology, knowledge, and context, the ways these factors impact and are obtained vary greatly. The current division of agents is not detailed enough, each agent’s design is overly simplistic, and the information aggregation process is quite arbitrary, failing to fully exploit the potential of LLMs.
6 Benchmark & Evaluation
This section compiles various benchmarks, evaluation methods, and metrics for Affective Computing, encompassing both Affective Understanding and Affective Generation. Table II comprehensively summarizes these benchmarks, detailing the evaluated tasks, datasets, metrics, and models.
TABLE II: Various benchmarks for Affective Computing. The AE means Automatic Evaluation, the HE means Human Evaluation.
Name Tasks Datasets Metrics Models
Text Annotation [96] Sentiment Analysis Sports tweets and Amazon reviews AE (Accuracy, F1, …) GPT-3.5-turbo
SOUL [114] Sentiment Analysis Yelp [182], IMDb [183] AE (F1, BLEU, …), HE Flan-T5, GPT-3.5-turbo
GPT-4V with Emotion[184] Sentiment Analysis MVSA-Single, DFEW, etc. 21 datasets AE (Accuracy, WAR, WAF) GPT-4V
MERBench [185] Multimodal Sentiment Analysis MER2023, etc. 7 datasets AE (WAF, MSE) PLMs, LLMs
MC-EIU [84] Emotion and Intent Joint Understanding in Multimodal Conversation MC-EIU test AE (WAF) EI2, other PLMs
LLMs2ESC [174] Emotional Support Conversation ESConv AE (Proficiency, Preference bias, ..) , HE GPT series, other LLMs
Cue-CoT [61] Emotional Support Conversation ED, etc. 6 datasets AE (Avg BLEU and F1), HE ChatGLM, other LLMs
ECoT [177] Emotional Support Conversation IEMCOP, COCO, et 7 datasets AE (EGS) GPT series, other LLMs
FEEL [186] Emotional Support Conversation ESCEval AE (Spear., Kend.), HE : GPT-3.5-turbo, ERNIE-4.0, GLM-4
MEDIC [187] Empathetic Response Generation MEDIC test AE (EE, ER, and CR) TFN, other PLMs
EmotionBench [188] Empathetic Response Generation EmotionBench test AE, HE GPT series, other LLMs
ChatGPT2ERG [189] Empathetic Response Generation EMPATHETICDIALOGUES AE (Distinct-n, BERTscore, BLEU-n), HE GPT-3.5-turbo, other PLMs
STICKERCONV [55] Multimodal Empathetic Response Generation STICKERCONV AE(MMr, BLEU, …), HE PEGS, ChatGLM3, Vicuna
SECEU [190] Emotional Intelligence Evaluation SECEU test AE (SECEU score, EQ score, …) GPT series, and other LLMs
EQ-Bench [113] Emotional Intelligence Evaluation EQ-Bench test AE (EQ-Bench scores) GPT series, other LLMs
EmoBench [191] Emotional Intelligence Evaluation EmoBench test AE (EmoBench scores) GPT series, other LLMs
EIBENCH [145] Emotional Intelligence, General Intelligence MMLU, etc. 88 datasets AE (Accuracy) Flan-T5, LLaMA-2
ChatGPT2AC [99] Affective Computing res14, et 13 datasets AE (Accuracy, UAR) gpt-4
MM-INSTRUCTEVAL [192] Multimodal Reasoning with Multimodal Contexts MVSA-Single etc. 16 datasets AE (Accuracy) GPT series, other LLMs
6.1 Affective Understanding
This section introduces the benchmarks in the field of Affective Understanding for different datasets related to SA.
Text Annotation [96] evaluates GPT-3.5-turbo as a tool for sentiment analysis labeling on two distinct datasets: sports tweets and Amazon reviews. GPT-3.5-turbo shows promise as a tool for text labeling for various sentiment analysis events and tasks.
Different from merely predicting the polarity of sentiment, a novel task called Sentiment and Opinion Understanding of Language (SOUL) is introduced, including two subtasks: Review Comprehension (RC) and Justification Generation (JG) [114]. RC seeks to validate statements that focus on subjective information based on a review text, while JG requires models to explain for their sentiment predictions.
GPT-4V with Emotion [184] presents the quantitative evaluation results of GPT-4V on 21 benchmark datasets covering six tasks: visual sentiment analysis, tweet sentiment analysis, micro-expression recognition, facial emotion recognition, dynamic facial emotion recognition, and multimodal emotion recognition. This paper collectively refers to these tasks as “Generalized Emotion Recognition (GER)”.
MERBench [185] is a benchmark for Multimodal Sentiment Analysis, covering of seven datasets: IEMOCAP [193], MELD [131], CMU-MOSI [194], CMU-MOSEI [195], CH-SIMS [196], CH-SIMS V2 [197]. MER2023, consisting of 3 datasets: MER-MULTI, MER-NOISE, MER-SEMI, is proposed as a Chinese benchmark. MERBench uses Automatic Evaluation, which includes the weighted average F-score (WAF) and the mean squared error (MSE).
MC-EIU dataset [84] is the first comprehensive and rich multimodal conversational emotion and intent joint understanding dataset, consisting of 4,970 conversational video clips from TV series in Chinese and English. It uses Automatic Evaluation, the Weighted Average F-score (WAF), as the metric. Emotion and Intent Joint Understanding in Multimodal Conversation (MC-EIU) aims to decode the semantic information manifested in a multimodal conversational history while simultaneously inferring the emotions and intents for the current utterance.
6.2 Affective Generation
This section introduces the benchmarks in the field of AG covering Emotional Support Conversation and Empathetic Response Generation.
6.2.1 Emotional support conversation
Emotional support conversation (ESC) aims to alleviate individuals’ emotional intensity and provide guidance for navigating personal challenges through engaging dialogue. The performance of LLMs on ESConv is initially analyzed, called LLMs2ESC [174]. LLMs2ESC finds that wide range of LLMs exhibits different preferences for support strategies and proposes a new suite of metrics that focus on strategies: proficiency, preference, and preference bias.
Cue-CoT [61] is a benchmark consisting of 6 in-depth dialogue datasets in both Chinese and English, considering three major aspects of user statuses: personality, emotions, and psychology, exhibited during the conversation, forming a comprehensive evaluation benchmark incorporating various user statuses in the context of dialogue response generation. It uses Automatic Evaluation and Human Evaluation.
Emotional Chain-of-Thought (ECoT) is a plug-and-play prompting method that enhances the performance of LLMs on various emotional generation tasks by aligning with human emotional intelligence guidelines [177]. Emotional Generation Score (EGS) is an automatic evaluation method to evaluate the quality of LLMs in emotional conditioned generation on a benchmark. This benchmark contains both textual and multimodal subsets. The textual benchmark contains five datasets: IEMOCAP [193], DailyDialog [65], EmpatheticDialogues [198], ESConv [12], and PENS [199]. The multimodal benchmark contains two datasets, SentiCap [200] and COCO [201].
FEEL [186] is a framework for evaluating Emotional Support Capability (ESC) using LLMs as evaluators. It constructs ESCEval, a high-quality dataset composed of human evaluations of emotional support capabilities, using instances from ESConv [12] and AUGESC [202], and trains the weights of LLMs on this dataset. The evaluation of ESC is categorized into two dimensions: Emotional Support Skills (Informativeness, Comprehensibility, Helpfulness) and Text Quality (Consistency, Coherence, Safety), using various metrics.
6.2.2 Empathetic Response Generation
MEDIC [187] is the first multimodal empathy dataset for psychotherapy scenarios considering visual, audio, and textual modalities. This paper proposes a new conceptual method for describing empathy, consisting of three mechanisms: Expression of Experience (EE), Emotional Reaction (ER), and Cognitive Reaction (CR).
To evaluate the empathy ability of LLMs, etc., how their feelings change when presented with specific situations. EmotionBench [188] collects 428 distinct situations, categorized into 36 factors, and conducts a comprehensive survey in the emotion appraisal theory. EmotionBench sets up a direction to align the emotional responses of LLMs with humans, focusing on eight negative emotions.
A comprehensive empirical investigation is conducted on the performance of LLMs, represented by gpt-3.5-turbo, in empathetic dialogue. For convenience, we refer to it as ChatGPT2EGR [189]. This study proposes three improvement methods: semantically similar in-context learning, two-stage interactive generation, and combination with a knowledge base. It adopts Distinct-n (Dist-1/2), BERTscore (PBERT, RBERT, FBERT), and BLEU-n (B2/4) as the main automatic metrics for evaluating response generation performance and samples 100 dialogues from the testing dataset for human evaluation.
STICKERCONV [55] is a multimodal empathetic dialogue dataset, comprising users with 2K generated personalities, 12.9K dialogue sessions, and 5.8K unique stickers. Unlike other benchmarks, STICKERCONV generates textual empathetic responses and outputs stickers as a visual modality to enhance empathy. It uses Automatic Evaluation, Human Evaluation, and LLM-based Evaluation.
6.2.3 Emotional Intelligence
In the context of large language models (LLMs), emotional intelligence (EI) has emerged as a critical area of focus. EI encompasses the capacity to identify, comprehend, regulate, and apply emotions across diverse situations [203]. To this end, a standardized EI test, Situational Evaluation of Complex Emotional Understanding (SECEU), suitable for both humans and LLMs is developed [190].
A novel benchmark, EQ-Bench [113], is designed to evaluate aspects of EI in Large Language Models (LLMs). EQ-Bench test dataset is constructed by GPT-4 to generate the dialogues to serve as the context for the test questions.
EmoBench [191] is a theory-based comprehensive EI benchmark for LLMs evaluation, consisting of a set of 400 hand-crafted questions, available in English and Chinese. The framework presents an extensive definition for machine’s EI, covering its essential capabilities: emotional understanding and emotional application.
EIBENCH [145] is a large curated collection of EI-related tasks converted into a text-to-text format, covering 15 tasks with 88 datasets. It primarily focuses on three aspects of the EI domain: Emotion Perception, Emotion Cognition, and Emotion Expression. To evaluate a model’s General Intelligence, EIBENCH conducts evaluations across four crucial dimensions: World Knowledge, General Reasoning, Commonsense Reasoning, and Reading Comprehension.
6.3 Others
This section introduces other benchmarks that do not easily fit into AU or AG tasks. For instance, one benchmark evaluates AU and AG, while another assesses multiple multimodal reasoning tasks with multimodal contexts, including multimodal sentiment analysis, sarcasm detection, and other tasks.
Executing a wide evaluation of the performance of InstructGPT and GPT-4, on a wide range of affective computing tasks, including aspect extraction, aspect polarity classification, opinion extraction, sentiment analysis, sentiment intensity detection, suicide tendency detection, toxicity detection, well-being assessment, engagement measurement, personality assessment and sarcasm detection, etc.. For convenience, we call this evaluation as ChatGPT2AC [99]. The automatic evaluation metrics are Unweighted Average Recall (UAR) and Accuracy.
MM-INSTRUCTEVAL [192] conducts comprehensive evaluations and establishs benchmarks for 31 models, including 23 MLLMs, using ten instructions across 16 diverse multimodal reasoning datasets with vision-text contexts. It uses automatic evaluation metrics such as Best Performance, Mean Relative Gain, Stability and Adaptability.
7 Discussion and Future Works
  1. 1.
    Ethics
    Data is essential and invaluable in artificial intelligence. Sentiment data from the real world, such as data from social media, psychological counseling, is highly personal, sensitive, regional, and cultural. Therefore, ethics must be considered when constructing a dataset, including individual privacy protection, prejudice elimination, and value alignment. Acquiring, preprocessing, and annotating data are all very time-consuming and labor-intensive processes. Thus, leveraging the world knowledge and generative capabilities of LLMs to synthesize datasets for affective computing tasks is an important research direction. Furthermore, improving the quality of original datasets and balancing the new and old datasets is a problem worth studying.
  2. 2.
    Effective Evaluation
    Although LLMs have achieved excellent performance in various tasks, it remains essential to explore how to train an LLM that can effectively perceive sentiments and evaluate performance. When training LLMs dedicated to AC tasks, researchers face the challenge of not only retaining the original abilitie of modles, such as instruction understanding and reasoning, but also strengthening new abilities, specifically sentiment perception and cognition. Additionally, it is crucial to ensure the integrity of values, facts, and logic within the LLMs. In terms of evaluation, previous studies conducted before the emergence of LLMs proposed various metrics to measure model performance, such as Accuracy, BLEU, and ROUGE. However, these metrics are insufficient to fully assess the ability of LLMs, e.g., generalization, generation, perception, and cognitive. Establishing a comprehensive and effective evaluation standard remains an urgent problem to be solved.
  3. 3.
    Multilingual and Multicultural Affective Computing
    Multilingual and Multicultural sentiment analysis spans different languages and cultures, focusing on sentiment recognition and analysis. Users from different linguistic or cultural backgrounds express emotions differently; for example, the same body movements can have different meanings in various cultural contexts. Current sentiment analysis is typically conducted on a single language or culture. Advances in multicultural sentiment analysis will significantly enhance the understanding of global users’ emotional reactions and behavior patterns, thereby improving the quality of cross-cultural communication and business decision-making. When training the current LLMs, the quality and quantity of corpora in different languages vary greatly, resulting in uneven AC performance in different languages. The core challenge in this field is effectively processing and integrating information from multiple languages and cultures to achieve accurate sentiment analysis. Therefore, it is necessary to continue to improve the multilingual and multicultural capabilities of the LLMs from multiple perspectives.
  4. 4.
    Multimodal Affective Computing
    Large Language Models (LLMs) that focus solely on the text modality have been expanded to Multimodal Large Language Models (MLLMs) that can process both image and text modalities. In the field of AC, data from different modalities such as text, visual, audio, electroencephalogram (EEG), and electrooculogram (EOG) can express individual sentiments. Effectively collecting and utilizing this multimodal data to train unified MLLMs for affective computing tasks is crucial. Addressing challenges such as micro-expression recognition, audio/video sentiment capture, and advancing brain-computer interface and emotional intelligence technologies represents significant research directions.
  5. 5.
    Real-time Affective Computing
    The first step in sentiment analysis is building a dataset and performing sentiment analysis. Real-time sentiment analysis in a dynamic environment, by integrating data in real time and identifying and analyzing sentiment, is particularly critical for various practical applications such as online customer service, real-time monitoring, and interactive entertainment. Real-time sentiment analysis requires collecting and processing data in real-time, which presents several challenges—ensuring the real-time performance of data processing and sentiment analysis, especially maintaining time synchronization and consistency across different modalities. Real-time systems often encounter various environmental noises. Another major challenge is ensuring that the system can adapt to environmental changes and noise interference while maintaining the accuracy and stability of sentiment recognition.
  6. 6.
    Vertical Domain of Affective Computing LLMs
    LLMs perform well on most tasks. However, there is room for improvement in more challenging tasks, such as Joint Aspect-Based Sentiment Analysis and Emotion-Cause Pair Extraction. Models specifically tailored for affective computing within particular industry sectors or areas of expertise, such as depression detection, psychological counseling, and sentiment companionship, remain rare. Developing effective Vertical Domain LLMs for Affective Computing, capable of simultaneously handling both AU and AG tasks, is a valuable area of study.
8 Conclusion
Affective Computing (AC) has emerged as a crucial research direction in artificial intelligence. With the rise of Large Language Models (LLMs), significant progress has been made in understanding and generating emotions. This paper provides a comprehensive review of LLM applications in AC, exploring the roles of techniques such as instruction tuning and prompt engineering in affective understanding and generation tasks. We further conclude benchmarks for multiple large language models in AC to offer in-depth assistance to researchers and practitioners in related fields. Research indicates that LLMs, leveraging their powerful contextual learning and sequence generation capabilities, have demonstrated excellent affective understanding and generation performance. However, LLMs still face numerous challenges in multilingual and multicultural sentiment analysis, real-time affective computing, and emotion-related tasks in specific vertical domains. We discuss challenges and future research directions related to AC, and aim to provide insights into this rapidly advancing field.
Acknowledgments
This research is supported by the National Natural Science Foundation of China (No.62272092, No.62172086, and No.62106039).