### v2

Abstract
Affective Computing (AC) integrates computer science, psychology, and cognitive science to enable machines to recognize, interpret, and simulate human emotions across domains such as social media, finance, healthcare, and education. AC commonly centers on two task families: Affective Understanding (AU) and Affective Generation (AG). While fine-tuned pre-trained language models (PLMs) have achieved solid AU performance, they often generalize poorly across tasks and remain limited for AG, especially in producing diverse, emotionally appropriate responses. The advent of Large Language Models (LLMs) (e.g., ChatGPT and LLaMA) has catalyzed a paradigm shift by offering in-context learning, broader world knowledge, and stronger sequence generation. This survey presents an NLP-oriented overview of AC in the LLM era. We (i) consolidate traditional AC tasks and preliminary LLM-based studies; (ii) review adaptation techniques that improve AU/AG, including Instruction Tuning (full and parameter-efficient methods such as LoRA, P-/Prompt-Tuning), Prompt Engineering (zero/few-shot, chain-of-thought, agent-based prompting), and Reinforcement Learning. For the latter, we summarize RL from human preferences (RLHF), verifiable/programmatic rewards (RLVR), and AI feedback (RLAIF), which provide preference- or rule-grounded optimization signals that can help steer AU/AG toward empathy, safety, and planning, achieving finer-grained or multi-objective control. To assess progress, we compile benchmarks and evaluation practices for both AU and AG. We also discuss open challenges—from ethics, data quality, and safety to robust evaluation and resource efficiency—and outline research directions. We hope this survey clarifies the landscape and offers practical guidance for building affect-aware, reliable, and responsible LLM systems.
keywords:
[…]
1 Introduction
Affective Computing (AC), encompassing the study and development of systems capable of recognizing, interpreting, processing, and simulating human affects through natural language and other modalities, seeks to bridge the gap between human emotions and machine understanding Picard (2000); Tao & Tan (2005); Banafa (2024). The AC task can be broadly categorized into two mainstream tasks: the Affective Understanding (AU) task Zhao et al. (2023a) and the Affective Generation (AG) task Nie & Zhan (2022); Pei et al. (2024). Affective Understanding focuses on recognizing and interpreting human emotions, covering Sentiment Analysis Medhat et al. (2014); Taboada (2016), Sarcasm Detection Joshi et al. (2017); Verma et al. (2021), and so on. Affective Generation involves creating responses or content that can elicit specific emotional responses in humans to generate expressions with emotionally nuanced, including Emotional-aware Dialogue Generation Rashkin et al. (2019); Liu et al. (2021), Creative Content Generation, etc. With the emergence of Large Language Models (LLMs) OpenAI (2023); Touvron et al. (2023); Zeng et al. (2023), the Affective Computing task is experiencing a paradigm shift to explore and model human emotions in unprecedented ways.
Figure 1: Comparison of approaches to AC tasks across the PLMs and LLMs eras. In the PLMs era, task-specific datasets fine-tune PLMs with sentiment-aware modules and downstream DNNs Bengio et al. (2009) (e.g., MLP Rosenblatt (1958), LSTM Hochreiter & Schmidhuber (1997)), typically requiring a separate model per task. In the LLMs era, a single (mostly frozen) LLM handles multiple tasks via (i) instruction tuning with adapters, (ii) prompt engineering, and (iii) reinforcement learning that samples rollouts and uses reward/preference signals to optimize a loss and update the adapters.
Prior to the advent of LLMs, the dominant paradigm of Affective Computing involves fine-tuning Pre-trained Language Models (PLMs) Amin et al. (2023a), such as BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), on labelled datasets for specific tasks to improve performance. While the above training paradigm has achieved numerous successes in the tasks related to Affective Understanding, it still is inherently limited by the quality and scale of manually annotated data and struggles to generalize to new domains Mao et al. (2023). The rise of LLMs research, as demonstrated in Figure 1, profoundly transforms the research paradigm of AC, expanding task variety and surpassing the constraints of previous studies, which often focus on particular datasets and single-task the performance improvements Tan et al. (2023a); Krugmann & Hartmann (2024). This evolution is accompanied by developing novel approaches that exploit the vast knowledge of LLMs Wei et al. (2022a) to simultaneously boost the performance of various tasks related to Affective Computing.
The auto-regressive nature and massive pre-training of LLMs give them an advantage in processing dialogue inputs and generating coherent responses Radford et al. (2018); Brown et al. (2020). Various techniques are proposed to fully harness and enhance the capabilities of LLMs, including Instruction Tuning, Prompt Engineering and Reinforcement Learning. Specifically, Instruction Tuning Chung et al. (2024); Zhang et al. (2024d) involves fine-tuning a pre-trained language model on the dataset where tasks are explicitly defined through instructions. For example, Flan-T5 Chung et al. (2024) designs specific instructions for different tasks, e.g., Sentiment Analysis, Commonsense Reasoning, Question Answering, etc. Instruction Tuning is a form of targeted training that aims to make the model more adept at following complex and varied instructions across different domains. Prompt Engineering Zhou et al. (2023b); White et al. (2023) is more about crafting the input given to the model to elicit the desired output without modifying the model itself. In-context Learning can further mine a model to learn and apply knowledge based on the instances provided directly within the prompt, without any prior explicit training on these specific tasks Xie et al. (2022). Reinforcement Learning optimizes the model against preference-based or verifiable rewards to improve helpfulness, safety, and style. A common pipeline, Reinforcement Learning from Human Feedback (RLHF), first trains a reward model from human comparisons and then updates the policy with online RL—typically Proximal Policy Optimization (PPO) Lambert (2025). Complementary directions include Reinforcement Learning with Verifiable Rewards (RLVR) Lambert et al. (2024), which uses rule-checkers or pass/fail signals as rewards, and methods that reduce or avoid online RL, such as Direct Preference Optimization (DPO) Rafailov et al. (2023a). Reinforcement Learning from AI Feedback (RLAIF) distills feedback from capable AI evaluators or rule sets into the preference signal (e.g., Wen et al. (2025)). For AC tasks, these approaches allow alignment directly toward affect-aware objectives, covering politeness, empathy, and non-toxic style.
Although LLMs offer new opportunities and show initial success in various tasks, a comprehensive review that consolidates and summarizes the recent advancements in AC with LLMs from an NLP perspective still needs to be present. Existing surveys either focus narrowly on specific tasks like sentiment analysis or a limited set of classic affective computing tasks Tan et al. (2023a); Wake et al. (2023); Cui et al. (2023); Zhang et al. (2023e); Zhu et al. (2024), or they concentrate on traditional, non-LLM-based methods Amin et al. (2023a); Afzal et al. (2023). To mitigate the gap, it is essential to provide a comprehensive overview (Figure 2) of how LLMs are being applied across the diverse spectrum of tasks related to AC, along with the unique opportunities and challenges they bring. Accordingly, we investigate the capabilities and limitations of LLMs in AC tasks, covering both Affective Understanding and Generation. While our focus is primarily on NLP, we also include text-centric multimodal affective computing tasks, recognizing the growing importance of integrating multiple modalities in AC. We begin by summarizing early studies that laid the groundwork for understanding the capabilities and limitations of LLMs in AC. Building on this, we explore three critical technologies for adapting LLMs to AC tasks: instruction tuning (§4), prompt engineering (§5) and reinforcement learning (§6). While tuning grounds the model in task formats and prompting controls behavior at inference, RL supplies preference- or programmatically verifiable objectives that close the optimization loop. To deliver a comprehensive assessment, we further summarize benchmarks and evaluation methods (§7) for systematically evaluating LLMs’ performance in AC, including the innovative use of LLMs as evaluation tools. Finally, we enumerate open challenges and outline promising future research directions (§8), offering insights into this rapidly advancing field’s continuous evolution and potential.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center, text width=4cm, s sep=6pt, calign=child edge, calign child=(n_children()+1)/2, l sep=12pt, , [Affective Computing, root, [Preliminary Study (§3), preliminary [Affective Understanding (§3.1), preliminary [Tan et al. Tan et al. (2023b); Rodriguez et al. Rodr´ıguez-Ibánez et al. (2023); Das et al. Das & Singh (2023); Krugmann et al. Krugmann & Hartmann (2024); Zhang et al. Zhang et al. (2023f); Belal et al. Belal et al. (2023); Lossio et al. Lossio-Ventura et al. (2024); Alexander et al. Alexandersen & Rutlin (2024); Amin et al. Amin et al. (2023b); Wang et al. Wang et al. (2024h); Wu et al. Wu et al. (2024a); Zhang et al. Zhang et al. (2023d); M.M.Amin et al. Amin et al. (2023a), preliminary, text width=12cm ] ] [Affective Generation (§3.2), preliminary [ Zhao et al. Zhao et al. (2023a); Welivita et al. Welivita & Pu (2024); Elyoseph et al. Elyoseph et al. (2024); SECEU Wang et al. (2023e); EQ-Bench Paech (2024); SOUL Deng et al. (2023d), preliminary, text width=12cm ] ] ] [Instruction Tuning (§4), sft [Parameter-Efficient Fine-Tuning, sft [ EMO-LLaMA Xing et al. (2024); Emotion-LLaMA Cheng et al. (2024b); InstructERC Lei et al. (2024); CKERC Fu (2024); DialogueLLM Zhang et al. (2023h); Iain J. et. al Cruickshank & Ng (2024); STOEI Jia et al. (2024); EKTC Cao et al. (2025); Wang et al. Wang et al. (2023b); PEGS Zhang et al. (2024f); GSA Hou et al. (2024); Penf et al. Peng et al. (2024); Ding et al. Ding et al. (2024); M2SE Li et al. (2024a); MoEI Zhao et al. (2024b); Zhao et al. Zhao et al. (2024b); PICA Zhang et al. (2023i); Zheng et al. Zheng et al. (2023b); BLSP-Emo Wang et al. (2024a); EMOLLM Liu et al. (2024c); FAME-Net Wang et al. (2024e); PerceptiveAgent Yan et al. (2024), sft, text width=12cm] ] [Full-Parameter Fine-Tuning, sft [Colin Raffel et al. Raffel et al. (2020a); Boyu Zhang et al. Zhang et al. (2023a); USA Gan et al. (2023); Wisdom Wang et al. (2024f); Omni-Emotion Yang et al. (2025); InstructABSA Scaria et al. (2023); ITSCL Zhang et al. (2024b); Siddharth Varia et al. Varia et al. (2023); UniSA Li et al. (2023b); NUS-Emo Luo et al. (2024); ObG Wang et al. (2024b); SoulChat Chen et al. (2023b); EMOVA Chen et al. (2024a); Emollms Liu et al. (2024c); MODA Zhang et al. (2025c); Emoada Dong et al. (2024); Lim Dongjun et al. Lim & Cheong (2024), sft, text width=12cm] ] ] [Prompt Engineering (§5), prompt [Zero-shot (§5.1.1 & 5.2.1), prompt [ Amin et al. Amin et al. (2023b); Ouyang et al. Ouyang et al. (2024); M.A.Al Asad et al. Al Asad et al. (2023); W.Stigall et al. Stigall et al. (2024); Hong et al. Hong et al. (2025); Zhou et al. Zhou et al. (2023a); Bai et al. Bai et al. (2024); Cabello et al. Cabello & Akujuobi (2024); Wang et al. Wang & Luo (2023) Niu et al. Niu et al. (2024); Plaza et al. Plaza-del Arco et al. (2024); Li et al. Li et al. (2023a); Shen et al. Shen et al. (2024); Wu et al. Wu et al. (2024d); Santoso et al. Santoso et al. (2024); Liang et al. Liang et al. (2024); Wang et al. Wang et al. (2024g) Lee et al. Lee et al. (2023a); Yang et al. Yang et al. (2024c); TCG Bhaskar et al. (2022); Y.M.Resendiz Resendiz & Klinger (2023); PromptMind Su et al. (2023), prompt, text width=12cm]] [Few-shot (§5.1.2 & 5.2.2), prompt [ Zhou et al. Zhou et al. (2024); Zhang et al. Zhang et al. (2024d); Yang et al. Yang et al. (2024a); SCRAP Kim et al. (2024); Wisdom Yang et al. (2024a); LVLMS Liu et al. (2024a); Ye et al. (2024); SEGA Chen et al. (2024c); Sibyl Wang et al. (2023c); Chen et al. Chen et al. (2023a); Kang et al. Kang et al. (2024); LLaMA2 Touvron et al. (2023), prompt, text width=12cm]] [Chain-of-Thought (§5.1.3 & 5.2.3), prompt [ Lee et al. Lee et al. (2024); THOR Fei et al. (2023); Lyu et al. Lyu et al. (2024); CoE Lee et al. (2023b); Manzoor et al. Manzoor et al. (2024); ECR-Chain Huang et al. (2024); RP-CoT Wang & Luo (2023); DECC Wu et al. (2024b); EMO Baumann & Kramer (2024); Deng Deng et al. (2023a); CFEG Chen et al. (2024b); Cue-CoT Wang et al. (2023a); EDIT Wu et al. (2023); Aptness Hu et al. (2024); ProCoT Deng et al. (2023b); ECoT Li et al. (2024g); ESCoT Zhang et al. (2024e); Li et al. Li et al. (2024d); VEHICLE Shao et al. (2024a), prompt, text width=12cm]] [Agent-based (§5.1.4 & 5.2.4), prompt [ Fei et al. Fei et al. (2023); Yang et al. Yang et al. (2024a); Sun et al. Sun et al. (2023); Wei et al. Wei et al. (2024); E.S.Van et al. Van Haeringen et al. (2023); PANAS Regan et al. (2024); HAD Xing (2024); Agent4SC Zhang et al. (2024f); COOPER Cheng et al. (2024a); PerceptiveAgent Yan et al. (2024); Emotion-aware agents Hu et al. (2022b); Zhang et al. Zhang et al. (2024c), prompt, text width=12cm]] ] [Reinforcement Learning (§6), rl [RLHF(§6.1), rl [ Zhu et al. Zhu et al. (2023a); XU et al. Xu et al. (2025); AlignCap Liang et al. (2024); Chen et al. Chen et al. (2025), rl, text width=12cm] ] [RLVR(§6.2), rl [ R1-Omni Zhao et al. (2025a); SAGE (State Augmented GEneration) Zhang & Jaitly (2025); SAGE (Sentient Agent as a Judge) Zhang et al. (2025a); CRP Li et al. (2024c); EmpRL Ma et al. (2025); ReflectDiffu Yuan et al. (2024); ABLE Mishra et al. (2024); GENTEEL-NEGOTIATOR Priya et al. (2025); F2RL Wang et al. (2024d); Unnikrishnan et al. Unnikrishnan (2024), rl,text width=12cm] ] [RLAIF(§6.3), rl [ Yoshida et al. Yoshida et al. (2025) PPDPP Deng et al. (2023c); DialogXpert Rakib et al. (2025); CSO Zhao et al. (2025b); Ye et al. Ye et al. (2025); Decoupled ESC Zhang et al. (2025b), rl,text width=12cm] ] ] [Benchmark & Evaluation (§7), benchmark [Affective Understanding (§7.1), benchmark [SOUL Deng et al. (2023d); GPT-4v with Emotion Lian et al. (2023b); MERBench Lian et al. (2024); DFEW Jiang et al. (2020); MC-EIU Liu et al. (2024b), benchmark, text width=12cm] ] [Affective Generation (§7.2), benchmark [ Cue-CoT Wang et al. (2023a); ESC-Eval Zhao et al. (2024a); MEDIC Zhu et al. (2023b); EmotionBench tse Huang et al. (2023); STICKERCONV Zhang et al. (2024f); SECEU Wang et al. (2023d); EQ-Bench Paech (2024); EmoBench Sabour et al. (2024); EIBENCH Zhao et al. (2024b); EmoVIT Xie et al. (2024); EMER Lian et al. (2023a); MER-UniBench Lian et al. (2025), benchmark, text width=12cm] ] [Others (§7.3), benchmark [ ChatGPT2AC Amin et al. (2024); MM-INSTRUCTEVAL Yang et al. (2024b), benchmark, text width=12cm] ] ] ]
Figure 2: The overview of the paper.
2 Tasks
Affective Computing (AC) comprises two mainstream task families: Affective Understanding (AU) Zhao et al. (2023a) and Affective Generation (AG) Nie & Zhan (2022); Pei et al. (2024), as summarized in Table 1.
Table 1: The traditional tasks of Affective Computing from an NLP perspective, with representative methods in the LLM era.
AC Tasks Subtasks Examples
AU Sentiment Analysis (SA) Polarity Classification (PC) USA Gan et al. (2023), WisdoM Wang et al. (2024f), RP-CoT Wang & Luo (2023), Sun et.al Sun et al. (2023)
Emotional Classification (EC) Zhang et.al Zhang et al. (2023e), Stigall et.al Stigall et al. (2024)
Aspect-Based Sentiment Analysis (ABSA) InstructABSA Scaria et al. (2023), SCRAP Kim et al. (2024)
Emotional Intensity Detection (EID) Amin et.al Amin et al. (2023b)
Implicit Sentiment Analysis (ISA) THOR Fei et al. (2023)
Emotion Recognition in Conversation (ERC) UniMSE Hu et al. (2022a), DialogueLLM Zhang et al. (2023h), CKERC Fu (2024), InstructERC Lei et al. (2024)
Emotion Cause Paired Extraction (ECPE) DECC Wu et al. (2024b), ECR-Chain Huang et al. (2024)
Subjective Text Analysis (STA) Suicide Tendency Detection (STD) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Personality Assessment (PA) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Toxicity Detection (TD) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Sarcasm Detection (SD) Hoffmann et.al Hoffmann et al. (2022), Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Well-Being Assessment (WBA) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
Engagement Measurement (EM) Zhang et.al Zhang et al. (2023e), Amin et.al Amin et al. (2023b)
AG Emotional Dialogue (ED) Empathetic Response Generation (ERG) Sibyl Wang et al. (2023b), PromptMind Yang et al. (2024c), PEGS Zhang et al. (2024f), Chain of Empathy Lee et al. (2023b), ECoT Li et al. (2024f)
Emotional Support Conversation (ESC) SoulChat Chen et al. (2023b), EmoLLM EmoLLM (2024), PICA Zhang et al. (2023i), Cue-CoT Wang et al. (2023a)
Review Summarization (RS) Opinion Summarization (OS) TCG Bhaskar et al. (2022), LASS Li et al. (2023a)
2.1 Affective Understanding
2.1.1 Sentiment Analysis
AU focuses on recognizing and interpreting human affect, with Sentiment Analysis (SA) as the core task. The basic form is Polarity Classification (PC) Pang et al. (2002), which labels a text as negative, positive, or neutral. Emotional Classification (EC) Li et al. (2017); Demszky et al. (2020) extends this to finer-grained emotion categories. Aspect-Based Sentiment Analysis (ABSA) Schouten & Frasincar (2015); Zhang et al. (2023g) targets aspect-level opinions. Typical subtasks include Aspect Term Extraction (ATE) Toh & Wang (2014); Li et al. (2018) and Aspect-oriented Sentiment Classification (ASC) Lau et al. (2014, 2018); Joint Aspect–Sentiment Analysis (JASA) Zhuang et al. (2020) models both simultaneously, and ACOS Cai et al. (2021) further incorporates category and opinion dimensions. Beyond label granularity and aspects, Emotional Intensity Detection (EID) Bonnet et al. (2015); Alonso et al. (2015) estimates strength, Implicit Sentiment Analysis (ISA) Tubishat et al. (2018); Wei et al. (2020a) captures sentiment not explicitly expressed, and Multimodal Sentiment Analysis (MSA) Das & Singh (2023); Jung et al. (2022) fuses textual, visual, and acoustic cues.
Conversation-level understanding has become central with advances in human–computer interaction Chen et al. (2017). Emotion Recognition in Conversation (ERC) Li et al. (2017); Poria et al. (2019b, 2018) predicts utterance-level emotions while tracking dynamics in context. Emotion and Intent Joint Understanding in Multimodal Conversation (MC-EIU) jointly infers current-utterance emotion and intent Liu et al. (2024b). Emotion Cause Pair Extraction (ECPE) Xia & Ding (2019a) extracts emotions together with their causes. See Table 1 for the complete taxonomy and representative methods.
2.1.2 Subjective Text Analysis
Subjective Text Analysis (STA) probes user traits, risks, and attitudes expressed in text, including Personality Assessment Ponce-López et al. (2016), Suicide Tendency Detection Desu et al. (2022), Toxicity Detection [^1] https://kaggle.com/competitions/jigsaw-toxic-comment-classification-challenge, Sarcasm Detection Misra & Arora (2023), Well-being Assessment Rastogi et al. (2022), and Engagement Measurement [^2] https://www.kaggle.com/datasets/thedevastator/social-media-interactions-on-tedtalks-dataset. Compared with classic SA, these tasks are relatively under-explored and under-resourced.
2.2 Affective Generation
2.2.1 Emotional Dialogue
Affective Generation (AG) aims to produce emotionally appropriate, empathetic, or supportive content. In Emotional Dialogue, Empathetic Response Generation (ERG) Rashkin et al. (2019) generates context-aware empathetic replies, while Emotional Support Conversation (ESC) Liu et al. (2021) targets goal-directed support. Multimodal variants (e.g., MERG) leverage audio/vision with text to enhance empathy and naturalness Jung et al. (2022); Zhang et al. (2024f).
2.3 Review Summarization
Review Summarization (RS) is a broader task where Opinion Summarization (OS) Kim et al. (2011); Moussa et al. (2018) condenses many user opinions on a topic—often from social media or UGC—into concise summaries that surface overall sentiment and key points. Challenges arise from noisy, unstructured text and figurative or diverse language use (e.g., sarcasm).
It is worth noting that research on AG lagged behind AU in the pre-LLM era due to limitations in model capabilities and technology. However, with their exceptional sequence generation abilities, the advent of LLMs has ushered in new possibilities for affect generation, making it a focal point of research in the LLM era Park et al. (2023).
3 Preliminary Study
Large Language Models (LLMs), with their powerful natural language processing, are rapidly transforming the field of Affective Computing. This section explores the emerging applications of LLMs in crucial areas of Affective Computing: Affective Understanding and Affective Generation. We review the current state of research, highlighting the strengths and limitations of LLMs in these tasks and suggest promising directions for future exploration.
3.1 Affective Understanding
3.1.1 Sentiment Analysis
Sentiment Analysis (SA), a classic task in the Affective Computing domain, focus on the automatic identification and categorization of emotions and sentiments expressed through contents from users Tan et al. (2023b); Rodr´ıguez-Ibánez et al. (2023); Das & Singh (2023).
To evaluate the performance of LLMs in SA tasks, researchers conduct assessments on multiple datasets and discover that the performance of LLMs is closely related to the complexity of the SA tasks they face. For more straightforward sentiment analysis tasks, such as Polarity Classification (PC), closed-source large language models like GPT-3.5-turbo demonstrate robust zero-shot learning capabilities and can deliver satisfactory results Krugmann & Hartmann (2024); Zhang et al. (2023f); Belal et al. (2023); Lossio-Ventura et al. (2024); Alexandersen & Rutlin (2024). However, when confronted with complex SA tasks requiring deep semantic understanding or structured emotional information, such as Emotional Classification (EC), Aspected-Based Sentiment Analysis (ABSA), Emotional Recognition in Conversation (ERC), etc., which involves identifying multiple sentiments corresponding to different aspects of a post, or Emotion Cause Analysis, the performance of LLMs still lags behind that of pre-trained language models (PLMs) fine-tuned on specific datasets Zhang et al. (2023f); Amin et al. (2024); Wang et al. (2024h); Wu et al. (2024a). This discrepancy may stem from LLMs lacking training data tailored to specific tasks and having insufficient capabilities to process complex semantic information.
Although the zero-shot learning capabilities of open-source LLMs are inferior to those of closed-source models, their superior few-shot learning abilities give them an advantage on datasets with limited training data or imbalanced distributions, outperforming PLMs in such scenarios. However, when data is abundant and label distribution is balanced, PLMs can still perform better due to their fine-tuning advantage on specific datasets Zhang et al. (2023d). Notably, LLMs show superior generalization capabilities in SA tasks compared to PLMs. By simply altering the prompts, LLMs can adapt to different datasets and tasks without requiring model retraining Wang et al. (2024h). However, the performance of LLMs is also affected by prompts and decoding parameters. For example, minor text attacks like synonym substitution can lead to fluctuations in their performance. Additionally, because of the auto-regressive decoding mechanism, parameters such as temperature and Top-P can significantly impact the effectiveness of LLMs in sentiment analysis tasks. Typically, a lower temperature value (below 0.3) and a lower Top-P value are necessary for achieving better results Amin et al. (2023a); Amin & Schuller (2024). This suggests that although research shows carefully crafted prompts can greatly improve LLMs’performance, developing optimal prompts and setting effective hyper-parameters for specific tasks and contexts remains an unresolved research challenge.
3.1.2 Subjective Text Analysis
LLMs demonstrate considerable performance variations across various Subjective Text Analysis (STA) related tasks. For instance, they excel in tasks involving negative emotion recognition, such as well-being assessment and toxicity detection. The superior performance could be due to the focus on safety and value alignment during the Reinforcement Learning from Human Feedback (RLHF) stage of LLM training Zheng et al. (2023a); Amin et al. (2023a, 2024). However, in tasks requiring the understanding of implicit emotional cues, such as engagement measurement, personality assessment, and sarcasm detection, LLMs exhibit a performance gap compared to PLMs fine-tuned on task-specific datasets. This discrepancy may stem from a need for more specialized training data during the supervised fine-tuning (SFT) phase of LLM development. Optimizing prompts for LLMs can enhance their performance on these specific tasks Amin et al. (2023a, 2024). Furthermore, leveraging their robust capacity for capturing contextual information, LLMs consistently demonstrate superior performance in processing longer texts compared to shorter texts across various subjective text analysis tasks Amin et al. (2024); Krugmann & Hartmann (2024). LLMs perform slightly worse than SFT models tailored to specific tasks in complex sentiment computing areas like sarcasm detection and subtle sentiment analysis. However, the substantial computational and time resources needed to fine-tune these large language models present a significant hurdle. Consequently, developing LLMs with outperformance and minimal training for specific tasks is an important research priority.
3.2 Affective Generation
3.2.1 Emotional Dialogue
Empathetic Response Generation (ERG) and Emotional Support Conversation (ESC) are two prominent tasks in AG that attract significant attention, utilizing the EMPATHETICDIALOGUES Rashkin et al. (2019) and ESConv Liu et al. (2021), respectively. Early research suggests that LLMs possess immense potential in emotionally-aware generation tasks. For example, GPT-3.5-turbo has shown competitive performance across multiple datasets Zhao et al. (2023a), while GPT-4 generates empathetic responses in various emotional scenarios. Human evaluations indicate that GPT-4’s responses exceed human performance by an average of 10% in empathy levels Welivita & Pu (2024). Unlike earlier models, LLMs produce longer and more diverse text Welivita & Pu (2024), introducing new evaluation challenges. Traditional automatic evaluation methods based on overlap metrics need to adequately measure the performance of LLMs in emotional generation tasks, while manual evaluation is costly and burdensome to scale. Developing novel evaluation methods and establishing standardized frameworks to accurately measure LLM-generated affective responses’ quality, empathy, and emotional intelligence is crucial for advancing research in this area.
3.2.2 Emotional Intelligence
Emotional Intelligence (EI) Dulewicz & Higgs (2000); Mayer et al. (2003); Neubauer & Freudenthaler (2005) is the ability to recognize, understand, manage, and use emotions effectively in various contexts. EI is crucial for enhancing the naturalness and effectiveness of human-computer interaction and is a vital component in achieving Artificial General Intelligence (AGI). To explore whether LLMs possess EI, researchers undertake investigations from various perspectives. Current studies Elyoseph et al. (2024); Ratican & Hutson (2023) draw upon assessment methods from the field of human psychology, employing tools, like the Levels of Emotional Awareness Scale (LEAS) Lane et al. (1990), to evaluate the ability of LLMs to perceive emotions. LLMs can generate relatively appropriate emotional responses, which has been evidenced.
However, using human-centric evaluation systems to measure the EI of LLMs has limitations, including the fact that the scales are designed with humans in mind, not explicitly specifically for LLMs and that the evaluations rely on human experts to conduct the assessments. Researchers have developed evaluation tools designed explicitly for LLMs to address these issue, such as the Situational Evaluation of Complex Emotional Understanding (SECEU) Wang et al. (2023e). Moreover, to improve the objectivity and efficiency of evaluations, EQ-bench has been proposed Paech (2024). It removes the need for expert assessments and instead directly evaluates the ability of LLMs to understand complex emotions and social interactions by predicting the intensity of emotional states in dialogue characters. While LLMs have demonstrated remarkable performance on SECEU and EQBench, which focus more on the cognition and understanding of human emotional scenarios, there is still insufficient evidence to suggest that they possess emotional intelligence comparable to humans. To address this gap, researchers propose a new task called Sentiment and Opinion Understanding of Language (SOUL), which aims to comprehensively evaluate the model’s ability to understand complex emotions through two sub-tasks, including Review Comprehension and Justification Generation. However, experimental results reveal a substantial performance gap of up to 27% between LLMs and human performance on the SOUL Deng et al. (2023d). It highlights the considerable room for improvement in the capacity of LLMs to understand the nuances and complexities of human emotions.
4 Instruction tuning
LLMs show excellent performance on different tasks, such as Question Answering, Sequence Generation, and so on Zhao et al. (2023b). Instruction Tuning Zhang et al. (2023c) is introduced to adapt LLMs to specific tasks, thereby maintaining the benefits of the original model while improving the performance of LLMs on specialized downstream tasks. Due to the substantial size of LLMs, which often exceed ten billion parameters, Full Parameter Fine-Tuning (FPFT) of all parameters of the complete model could be more practical. To reduce the cost of model training, a series of Parameter-Efficient Fine-Tuning (PEFT) approaches are proposed, such as LoRA Devalal & Karthikeyan (2018), Prefix Tuning Li & Liang (2021), and P-Tuning Liu et al. (2022), as shown in Figure 3. LoRA reduces the number of parameters that need to be trained during the fine-tuning process by freezing all the original parameters and injecting a pair of low-rank decomposition matrices next to the original weights. Prefix Tuning adds trainable "prefix" tokens related to specific tasks as an implicit learnable prompt to the input sequence of the LLM, inserting trainable free vectors into each layer of the Transformer. Applying PEFT in instruction tuning can effectively reduce the training costs in AC tasks.
(a) Full-Model Tuning
(b) Multi-LoRA
(c) LoRA
(d) Prompt Tuning
Figure 3: Illustrations of the instruction tuning methods. Pink blocks denote trainable modules, blue blocks signify frozen modules, and orange blocks represent optionally added modules. Subfigure 3(a) depicts the FPFT method, the transformer blocks of two distinct model architectures are fine-tuned. Subfigure 3(b) introduces the multi-LoRA method. In this approach, only the added multi-LoRA models are trained. This method allows for the optional incorporation of router modules to manage the weights of each LoRA module. Subfigure 3(c) demonstrates the LoRA method, The traditional LoRA method involves the addition of only two low-rank matrix modules, LoRA_A and LoRA_B. VeRA Kopiczko et al. (2024) allows the addition of two trainable vectors, d and b, augmenting the LoRA layer. This approach also shares LoRA_A and LoRA_B’s parameters to every LoRA layer. Moreover, when the rank of LoRA in each layer varies, this method can transition into adaLoRA Zhang et al. (2023b). Subfigure 3(d) illustrates the soft prompt tuning, where a trainable prefix is appended to the model’s embedding layer, known as p-tuning. Furthermore, scientific prefixes can be optionally added to each layer of the model, with only the added prefixes being trained, a technique referred to as prefix-tuning.
4.1 Affective Understanding
In addressing affective computing tasks, PLMs typically necessitate separate approaches for affective understanding and generation tasks. In contrast, LLMs can unify all affective computing tasks into a sequence generation format. However, directly employing LLMs for AU tasks does not offer significant advantages over fine-tuned PLMs. Consequently, numerous recent studies focus on constructing task-specific instruction tuning templates tailored to the unique characteristics of various tasks. These studies apply instruction tuning to the models, aiming to enhance the performance of LLMs on specific tasks or within particular domains.
4.1.1 Polarity and Emotional Classification
Polarity Classification (PC) and Emotional Classification (EC) tasks are frequently employed to assess a model’s comprehension and perceptual abilities. Several studies pioneer the exploration of the effects of instruction tuning of LLMs on the SA task Raffel et al. (2020b); Peng et al. (2024). The T5 Raffel et al. (2020a) series model, in particular, demonstrates an improved understanding of the sentiment content of text after instruction tuning. Recent research Peng et al. (2024) focuses on exploring the different PEFT methods, adopting the LoRA and P-tuning methods to fine-tuning ChatGLM2 GLM et al. (2024). Experimental results demonstrate that after instruction tuning, LLMs are greatly improved in eight datasets compared with the model without instruction tuning. However, the experimental results must conclusively determine which instruction tuning method is more effective in enhancing performance on PC.
While these studies demonstrate LLMs’ effectiveness in PC and EC tasks, it is equally important to understand their underlying sentiment processing mechanisms. Research shown that words with sentiment information can influence the overall sentiment of the text Gan et al. (2023). Building on this understanding, LLMs leverage techniques of instruction tuning to grasp word polarity, enabling them to effectively perform methods PC and EC tasks. These studies primarily focus on text-based PC and EC tasks, it is essential to recognize that real-world user expressions are not limited to text alone. User-generated content often encompasses multiple modalities, including images and audio, which can provide additional context and nuance to sentiment analysis. WisdoM Wang et al. (2024f) leverages contextual world knowledge derived from MLLM to enhance multimodal sentiment analysis (MSA). Additionally, the authors introduce a training-free module called "contextual fusion", which can reduce noise in the contextual data and improve model classification accuracy. WisdoM significantly outperforms existing state-of-the-art methods on the MSED dataset, demonstrating its effectiveness in integrating contextual knowledge to improve emotion classification. In WisdoM, the method of converting images into text captions results in the loss of a significant amount of image information and fails to focus on the more detailed emotional expressions within the images. EMO-LLaMA Xing et al. (2024) addresses this issue by training a dedicated encoder for facial features to extract both global and local information from faces. It employs handcrafted prompts to characterize user profiles and combines the previously extracted facial information to fine-tune the LLaMA model, thereby alleviating the model’s inaccuracies in recognizing image details. This model achieved the highest accuracy on both static and dynamic facial expression recognition datasets. Emotion-LLaMA Cheng et al. (2024b) additionally encodes audio; it uses a specific emotional encoder to seamlessly integrate audio, video, and text modalities, and performs LoRA fine-tuning on the LLaMA2 model. It achieved the highest score on the EMER dataset. Omni-Emotion Yang et al. (2025) has a similar model structure to Emotion-LLaMA and explores various fusion schemes of facial features with other visual features. Experiments found that the method of separately extracting the visual and facial features of the entire video and then concatenating them has the least impact on MLLMs.
4.1.2 Aspect-Based Sentiment Analysis
Aspect-Based Sentiment Analysis (ABSA) encompasses multiple subtasks, such as Aspect Term Extraction (ATE), Aspect-oriented Sentiment Classification (ASC), Joint Aspect-Sentiment Analysis (JASA), and Aspect-Category-Opinion-Sentiment (ACOS), and others. Traditional PLMs often design a separate module for each ABSA subtask, which reduces the correlation between ABSA subtasks. LLMs can leverage their strong generalization abilities to construct a unified framework for all ABSA subtasks.
Recent research explore this potential. InstructABSA introduces Instruction tuning based on the Tkk-Instruct model Wang et al. (2022) for ABSA. This study finds that LLMs are sensitive to different instructions, with experimental results fluctuating by up to 10% Scaria et al. (2023). ITSCL Zhang et al. (2024b) is based on the T5 model and performs unified encoding of emotions and opinions. It introduces contrastive learning loss at each layer of the model to differentiate the feature vectors of different emotions and entity labels, thereby enhancing the model’s performance. In addition to strengthening the correlations between ABSA tasks, large language models can leverage their extensive world knowledge to enhance information exchange across different domains. Some reseraches Ding et al. (2024) propose an ABSA cross-domain learning fine-tuning framework. In the first stage, domain knowledge is used to train the domain variation adapter, while replay data is used to train the domain-invariant adapter, employing orthogonal constraints to force the model to learn the differences between domain-invariant and domain-variant knowledge. In the second stage, all replay data will be used to train the domain-invariant adapter while freezing the corresponding domain variation adapter. This two-stage fine-tuning effectively utilizes knowledge from different domains and avoids the issue of forgetting previous domain knowledge. The model achieved the best performance across 19 datasets.
The ABSA task shares similar objectives with sentiment analysis and sentiment polarity analysis. Unisa aims to build a unified multimodal emotion analysis model that integrates various sentiment analysis tasks Li et al. (2023b). This model has already established a new dataset and benchmark, SAEval, to facilitate research in this field. SAEval encompasses various ABSA sub-tasks, providing a comprehensive evaluation platform for future research.
4.1.3 Emotion Recognition in Conversation
Emotion Recognition in Conversation (ERC) tasks involve recognizing the sentiment label of each utterance in the conversation, which is highly dependent on the conversation history and the speaker. InstructERC addresses these challenges by developing a template that includes a system prompt, dialogue history, and the label map. It also introduces Instruction Tuning using LoRA for ERC tasks Lei et al. (2024). Additionally, InstructERC creates two auxiliary tasks, speaker identification and emotion impact prediction, to enhance the model’s understanding of speaker information. This model performs very well on the MELD Poria et al. (2019a) and EmoryNLP Zahiri & Choi (2018) datasets.
Researchers extend ERC tasks to incorporate multimodal inputs, considering the frequently multimodal aspects of human communication, leading to the development of Multimodal Emotion Recognition in Conversation (MERC). DialogueLLM Zhang et al. (2023h) generates text descriptions of visual information through GPT-4, combines these descriptions with the conversation history as an instruction dataset, and uses LoRA to train LLaMA2. Through experiments, it has been found that adding information from other modalities can improve the model’s ability to understand dialogue information.
4.1.4 Emotion Cause Extraction
Emotion Cause Extraction (ECE) has garnered widespread attention in recent years. This task aims to extract potential causes behind certain emotions in text, but its application scenarios are significantly limited due to the need for annotating emotional information. To address this issue, the Emotion-Cause Pair Extraction (ECPE) task propose. The ECPE task simultaneously extracts potential Emotion-Cause pairs and their corresponding causes within a document, thereby reducing the dependence on emotion labels Xia & Ding (2019b). Most current research on ECPE tasks involves fine-tuning traditional PLMs Ding et al. (2020); Wei et al. (2020b) or leveraging the understanding capabilities of LLMs using well-designed prompts Wu et al. (2024b). These methods primarily focus on the Textual Emotion-Cause Pair Extraction in Conversations task (TECPE). However, emotion causes may also be embedded in other modalities, leading to the development of Multimodal Emotion-Cause Pair Extraction in Conversations task (MECPE) Wang et al. (2024c). LLMs shows potential in better integrating information from multiple modalities for such tasks.
Recent work on multimodal emotion-cause understanding has progressed from clause extraction to explicit cause generation. NUS-Emo Luo et al. (2024) attacks the original MECPE task with a two-stage self-training pipeline: LLMs are first emotion-tuned on ERC data, ImageBind aligns visual and textual affective representations, and the resulting multimodal predictions are recycled as pseudo-labels for further fine-tuning, yielding steady gains in cross-modal cause inference. Moving beyond extraction, Wang et al. introduce the Multimodal Emotion Cause Generation in Conversations (MECGC) task and the ObG framework Wang et al. (2024b); an emotion-cause-aware captioner (ECCap) is trained to verbalise cause-focused video descriptions, which are then fed to an encoder–decoder generator, demonstrating that caption-based bridging markedly improves reason generation. FAME-Net Wang et al. (2024e) refines this idea by fusing holistic scene features from a visual encoder with fine-grained facial cues from VGG16, allowing dual-granularity guidance that localises both contextual and expressive triggers. Finally, Emotion-LLaMA Cheng et al. (2024b) instruction-tunes LLaMA with a dedicated “[Reason]” token so that emotion prediction and textual justification are produced jointly, enhancing model confidence and interpretability. Collectively, these studies highlight the effectiveness of ERC pre-training, caption-based intermediaries, multi-stream visual fusion, and explanation-oriented instruction tuning for multimodal emotion-cause modelling.
4.1.5 Subjective Text Analysis
LLMs are employed to explore the Stance Classification task. In this task, the model determines whether a user agrees, disagrees, or is neutral on a given point of view Cruickshank & Ng (2024). The study investigates ten open-source models and seven prompting schemes, revealing exciting conclusions. These include the fact that LLMs are competitive with in-domain supervised models and that the fine-tuning process only sometimes leads to better performance than zero-shot models. This is likely because the fine-tuning makes the model too specialized, thus unable to generalize to out-of-domain data points and varying stance definitions. In the opinion expression task, STOEI Jia et al. (2024) concatenates text embeddings and speech embeddings, inputting them together into the LLM for instruction fine-tuning, enabling the model to learn to extract opinion words from the input and analyze the sentiments associated with those opinion words. This approach surpasses existing techniques by over 9.20% and achieves state-of-the-art (SOTA) results.
Although LLMs possess inherent advantages in these subjective sentiment analysis tasks, the datasets for these tasks are mostly scarce, making instruction fine-tuning of the model challenging.
4.2 Affective Generation
In the current landscape of artificial intelligence, content generation that expresses sentiment is applied to various practical fields, including emotional dialogue, chatbots, emotional support conversation, and more Nie & Zhan (2022). It enhances humanized services by analyzing and generating emotional information from users.
4.2.1 Emotional Dialogue
Emotional Dialogue (ED) refers to the ability to understand and share another person’s feelings and then respond in a way that conveys this understanding and compassion. LLMs must to possess high emotional intelligence to resonate with users in their replies. However, due to the limitations of the dataset and its relative homogeneity, the generated responses often tend to be formulaic and mechanical, resulting in lower interaction quality. Some researchers aim to construct a more diverse empathetic dialogue dataset. Considering the high labor costs associated with manually constructing datasets, some current studies utilize LLM-generated synthetic data to build high-quality empathetic dialogue datasets.Zheng et al. (2023b); Wang et al. (2023b); Zhang et al. (2024f).
To enlarge training resources for empathetic dialogue, several works exploit LLM-driven self-augmentation. SoulChat Chen et al. (2023b) iteratively prompts GPT-3.5-turbo to grow multi-turn exchanges, yielding richer emotional trajectories. EKTC Cao et al. (2025) constructs the ED-TooL dataset by asking LLaMA3 to label the exact moments at which external tools should be invoked; the resulting high-quality tool-call exemplars are then used for instruction-fine-tuning, enabling models to decide autonomously—and more cleanly—when tool usage will enhance their responses. In the task of empathetic responses, the information dimension carried by a single text modality is relatively limited. Stickers and voice can convey non-verbal cues in interpersonal communication, and multimodal interaction mechanisms hold significant value in enhancing empathetic expression. StickerConv generates a multi-scenario, multimodal empathetic dialogue dataset constructed from user information through a multi-agent system, enabling more natural and fluid interactions with users by generating stickers within the dialogue Zhang et al. (2024f).
Speech–centric approaches further exploit prosody to boost empathy. BLSP-Emo Wang et al. (2024a) follows a two-step alignment strategy: (i) convert the speaker’s audio into an explicit textual emotion hint and let an LLM draft a reply; (ii) apply QLoRA to inject the raw audio encodings themselves, forcing the model to ground its responses directly in vocal affect. PerceptiveAgent Yan et al. (2024) transcribes prosodic features with an audio-encoder–text-decoder pair, feeds the resulting description together with dialogue history to an LLM, and finally vocalises the generated reply via a synthesiser—thus closing the speech-to-speech loop. EMOVA Chen et al. (2024a) generalises this pipeline to multimodal inputs and, crucially, lets the model emit style tokens that steer the synthesiser, yielding richer and more engaging emotional speech.
4.2.2 Chatbot
Current datasets still limit most of the ERG and ESC tasks described above and need to capture the diversity of users and scenarios in real life fully. Chatbots capable of understanding user input and emotions are designed to address the practical application of emotional dialogue systems. These chatbots can offer encouragement during adversity and share in users’ happiness. PICA Zhang et al. (2023i) and EMOLLM EmoLLM (2024) are two notable examples of such chatbots. While LLMs can function well as "assistants," their emotional intelligence is often lacking, and long replies frequently fail to convey empathy. To address this issue, PICA is proposed. It uses P-tuning to fine-tune the ChatGLM2-6B model on two synthetic Chinese empathetic dialogue datasets, effectively enhancing the emotional capabilities of LLMs. EmoLLM is a series of large mental health models that can support the mental health counseling link of understanding users, supporting users, and helping users. It uses the QLoRADettmers et al. (2024) method to fine-tune multiple LLMs, such as Deepseek-R1_14b_int4, LLaMA3-8B-Instruct, and others, across different languages. By supporting various roles, EMOLLM enables the chatbot to generate more contextually appropriate user responses.
4.3 Multi-task of Affective Computing
Many experiments show that LLMs possess strong comprehension and generalization capabilities. Furthermore, their sequence generation ability enables them to construct a unified paradigm for AU and AG tasks. Consequently, researchers began exploring the development of a general AC model.
Due to the high similarity of mission objectives and approach in affective understanding tasks, researchers try to build a unified model that can accomplish tasks simultaneously. Emollms Liu et al. (2024c) develops a multi-task instruction dataset, AAID, covering SA and ERC. Experimental results show improved performance on each task after multi-task training. GSA-7B Hou et al. (2024) undergoes progressive LoRA training on three tasks, SA, ERC, and SD that are both related and increasingly challenging. It has been found that this hierarchical training approach can effectively improve the model’s performance across multiple tasks. In addition to classification tasks, M2SE Li et al. (2024a) also introduces tasks such as sentiment cause extraction, achieving a unified fine-tuning framework for multimodal and multitask scenarios. By sharing model parameters and facilitating knowledge transfer between tasks, it enhances the model’s ability to generalize in complex emotional contexts. This model is applicable to various emotion computation tasks.
However, current fine-tuning methods present limitations when dataset sources differ across tasks. Fine-tuning a model under these conditions can create conflicts and impact the original world knowledge embedded in the model Li et al. (2024b). These challenges highlight the need for more research on ensuring high-quality performance across all sentiment analysis tasks. To address these issues, researchers propose a novel Modular Emotional Intelligence enhancement method (MoEI) featuring two collaborative techniques—Modular Parameter Expansion (MPE) and Intra-Inter Modulation (I2M)—is proposed to fit most emotion analysis tasks Zhao et al. (2024b). In conjunction with this method, the study develops EIBENCH, a unified affective computing dataset containing 88 datasets across 15 EI-related tasks, is introduced. MoEI adopts the Mixture-of-LoRA (MoLoRA) method for fine-tuning to improve EI while maintaining the generative ability of the model.
In multimodality, MEILLMDong et al. (2024) integrates deep feature vectors extracted from visual, textual, and audio models into baichuan13B-chat through the MLP layer. After fine-tuning the instructions, he can complete a variety of emotional interaction tasks, including psychological assessment, psychological portrait, psychological counseling, etc. MODA Zhang et al. (2025c) develop a multimodal artificial intelligence that is both emotionally and intellectually capable. It addresses the issue of attention imbalance among different modalities in large models through a duplex attention alignment mechanism and a modular attention mask, allowing the model to focus more on fine-grained information from auxiliary modalities. In terms of cognitive and emotional understanding, MODA can accurately identify user intentions and emotional inclinations, demonstrating immense potential in the field of human-computer dialogue.
4.4 Challenge
While Instruction Tuning has enabled LLMs to make significant strides in AC tasks, several challenges remain to be addressed.
  • FPFT LLMs for specific domains are time-consuming, labor-intensive, and requires significant computational resources. While PEFT can reduce resource consumption in fine-tuning LLMs, they are more suitable for specific datasets and often sacrifice the generalization ability of LLMs. There is a need for scholarly research about the comparative efficacy of various PEFT methods on AC tasks. Consequently, developing a simple yet effective fine-tuning approach that balances efficiency, task-specific performance, and generalization capability remains a significant challenge in LLMs.
  • While LLMs with instruction tuning demonstrate excellent performance in AC tasks, they are still slightly inferior to the previous fully fine-tuned PLMs method in some tasks. Achieving state-of-the-art performance with minimal training remains a major challenge.
  • Despite their known universality, LLMs often struggle to simultaneously excel in both AU and AG within the domain of AC. It remains a significant challenge to develop a unified model capable of effectively addressing AU and AG.
  • Research in niche areas of AC tasks, such as financial sentiment analysis, sarcasm detection, and metaphor analysis, still needs to be improved. Furthermore, the full potential of LLMs in these specialized AC tasks has yet to be thoroughly explored and evaluated.
5 Prompt Engineering
Prompt engineering Brown et al. (2020) guides LLMs to produce desired outputs by designing appropriate prompts, effectively improving the accuracy and reliability of AC tasks. This section reviews recent advances in prompt engineering for AC, focusing on four techniques: zero-shot prompting, few-shot prompting, Chain-of-Thought (CoT) prompting, and agent.
Zero-shot Prompting Brown et al. (2020) generates text for a specific topic or domain without any demonstrations. It is simple to use, highly adaptable, and especially suitable for scenarios where data are scarce or hard to obtain.
Few-shot Prompting Brown et al. (2020) supplies the model with a small set of task-relevant examples or demonstrations. These examples, together with the task description, are used as prompts to help the model better understand the task, reason more effectively, and produce high-quality outputs.
Figure 4: The overview of prompt engineering.
Chain-of-Thought Wei et al. (2022b) adopts a multi-step reasoning paradigm that guides the LLMs to reason logically and explicitly reveal the intermediate steps of its reasoning chain. Chain-of-Thought enables LLMs to better leverage context, produce more accurate and interpretable answers, and perform better on complex reasoning tasks.
Agent Park et al. (2023) assigns clear roles, objectives, and collaboration protocols to multiple LLMs, driving them to collaborate via message passing and feedback loops to complete tasks. Agent effectively handles task decomposition and planning, tool invocation, and memory management, and excels in open-domain, multi-stage, highly interactive scenarios.
While these prompting paradigms share commonalities, they differ in emphasis and application contexts. Zero-shot prompting directs the model with task instructions alone, without examples. Few-shot prompting appends a small number of exemplars to the instruction to align format and decision boundaries. Chain-of-thought prompting explicitly exposes intermediate reasoning steps or uses trigger phrases to elicit step-by-step thinking. Agent methods orchestrate a workflow of one or more agents and tools to support task decomposition, planning, memory, and self-verification, and to aggregate multiple outputs. Figure 4 summarizes the distinctions and composability of the four approaches.
5.1 Affective Understanding
5.1.1 Zero-shot for Affective Understanding
In affective understanding (AU), zero-shot prompting has been applied. Researchers preset a task description as the prompt to the target text to guide LLMs to produce sentiment labels Amin et al. (2023b); Ouyang et al. (2024); Al Asad et al. (2023); Stigall et al. (2024). Building on this, incorporating dialogue context Hong et al. (2025), multi-label prompts Zhou et al. (2023a), constraint-based prompts Bai et al. (2024), and task-specific prefixes Cabello & Akujuobi (2024) further improves LLM performance on AU tasks. Researchers explore a variety of zero-shot prompting strategies to improve the performance of LLM-based models on the AU tasks. LLMs are cast as emotion experts Wang & Luo (2023), assigned specific personas Niu et al. (2024), or given gendered roles Plaza-del Arco et al. (2024) to perform sentiment analysis. Psychologically inspired prompts are also introduced, such as emotional stimulation questions Li et al. (2023a) and integrating narrative style elements extracted by HEART Shen et al. (2024). In Speech Emotion Recognition (SER), acoustic features are verbalized into textual descriptors Santoso et al. (2024), prompts are constructed in natural language Wu et al. (2024d), and emotional cues in speech are distilled into prompt prefixes Liang et al. (2024), thereby further improving recognition accuracy. Although fixed-template zero-shot prompting can be effective, its fixed content limits flexibility and diversity. Accordingly, multimodal sentiment analysis explores having LLMs generate task-specific prompts on demand and then using these prompts to guide the analysis of emotions in image–text pairs Wang et al. (2024f).
Additionally, subjective tasks such as Metaphor Recognition Mohler et al. (2016) and Dark Humor Detection Hoffmann et al. (2022) require greater attention to emotional responses. Reasoning in Conversation (RiC) Wang et al. (2024g) is a zero-shot paradigm that addresses tasks via dialogue simulation, extracting key information to support the final answer.
5.1.2 Few-shot for Affective Understanding
The pivotal role of demonstrations in prompt drisves exploration of the few-shot prompting. Early methods required manually curating and inserting demonstrations Zhou et al. (2024). With a range of selection strategies emerging Zhang et al. (2024d), a central challenge is how to more effectively elicit and unlock LLMs’ few-shot capabilities for sentiment analysis. For the challenging Aspect-Sentiment Quad Prediction (ASQP) task Zhang et al. (2021), The Self-Consistent Reasoning-based Aspect-Sentiment Quad Prediction model (SCRAP) Kim et al. (2024) integrates pre-generated reasoning demonstrations into the task’s reasoning process to strengthen LLM reasoning. Few-shot prompting has also been introduced into multimodal sentiment analysis (MSA) for AU. WisdoM Yang et al. (2024a) targets four multimodal entity-based sentiment analysis (MEBSA) subtasks by converting visual inputs into textual surrogates treated as auxiliary statements, and constructing demonstrations by selecting positive and negative instances based on textual similarity. During context generation, LVLMs Liu et al. (2024a); Ye et al. (2024) use prompt templates, together with the given image–sentence demonstrations, to produce contexts that explicitly incorporate world knowledge.
Additionally, in Depression Detection task, few-shot prompting guides data generation to enhancing the Structural Element Graph (SEGA) Chen et al. (2024c). Researchers follow principles to steer the synthesis process and provide the LLM with generation exemplars, yielding high-quality synthetic data.
5.1.3 Chain-of-Thought for Affective Understanding
Zero-shot or few-shot prompting carries limited information in a single-turn interaction, making it insufficient for complex AU tasks. To address this, researchers introduce CoT Lee et al. (2024), which explicitly adds intermediate reasoning steps to better handle complex tasks. Given the complexity of AU, THOR Fei et al. (2023) employs a three-hop CoT framework for Implicit Sentiment Analysis (ISA). It sequentially extracts fine-grained aspects, analyzes latent opinions, and infers implicit sentiment. In addition, injecting causal relations as causal cues into the model Lyu et al. (2024) can enhance LLMs’ sentiment analysis performance. Unlike traditional methods that extract regularities directly from text, Chain of Empathy (CoE) Lee et al. (2023b) combines multiple psychotherapeutic approaches Beck (1979); Linehan (1987); Cooper & McLeod (2011); Wubbolding et al. (2017), designs four types of chains of thought, and interprets users’ psychological states within a professional framework, thereby enabling a more precise understanding and handling of emotions. Meanwhile, other work Manzoor et al. (2024) adopts a psychological information–mining perspective and explores LLMs’ potential for empathetic understanding. ECR-Chain Huang et al. (2024) draws on appraisal theory’s “stimulus–appraisal–emotion” process and employs multi-step reasoning to infer the deeper causes of emotions.
Methods such as CoE and ECR-Chain generate task-specific reasoning and achieve good results, but they still face challenges on complex extraction tasks such as Joint Aspect Sentiment Analysis (JASA) and Emotion-Cause Pair Extraction (ECPE). Accordingly, RP-CoT Wang & Luo (2023) adopts multi-step reasoning to decompose the task: it first identifies aspect-level sentiment polarities and then determines the overall polarity. Likewise, DECC Wu et al. (2024b) first detects emotions and locates the corresponding emotion clauses, and then, for each emotion, selects the most probable cause clause and outputs emotion–cause pairs. Beyond generating intermediate reasoning for emotion understanding and task decomposition, CoT also supports prompt construction. The Evolutionary Multi-Objective (EMO) process Baumann & Kramer (2024) starts from human-written seed prompts, feeds them to an LLM and evaluates fitness via the generated stories. It then performs crossover recombination and applies mutation operators, iteratively producing new prompts.
The CoT can be combined with few-shot prompting. For Market Sentiment Analysis Chen et al. (2018), the complexity of financial and social-media terminology and data scarcity constrain traditional supervised methods. Researchers construct contextual exemplars and use few-shot prompts to elicit CoT summaries Deng et al. (2023a), prompting the model to retrieve and integrate domain knowledge before concluding and thereby injecting relevant financial knowledge into the Market Sentiment Analysis.
5.1.4 Agent for Affective Understanding
Prior studies mostly adopt a single LLM decision paradigm Fei et al. (2023); Yang et al. (2024a). However, a single model’s output does not reliably approach the optimum. To address this, Agent techniques orchestrate collaboration and evaluation among multiple LLMs, enabling more robust ensemble decisions.
To achieve multi-LLM negotiation in sentiment analysis, the Generator-Discriminator Role-switching Decision-Making framework Sun et al. (2023) is proposed. The generator injects reasoning and provides decisions with rationales, the discriminator explains and assesses credibility, and the two iterate until they reach consensus, leveraging model complementarity for more accurate emotion interpretation and correction. Building on this idea, agent design expands with an automated 5W1H question-generation module to predict emotions from social information Wei et al. (2024). Multi-agent interaction also builds an emotion contagion model that characterizes how emotions propagate within groups and their effects Van Haeringen et al. (2023). The PANAS framework Regan et al. (2024) employs three specialized agents for memory, context understanding, and evaluation. Meanwhile, AnnaAgent Wang et al. (2025a) uses multi-session memory to capture users’ long-term emotional dynamics, offering a practical route to more realistic and deeper digital companions.
In Financial Sentiment Analysis (FSA), HAD Xing (2024) summarizes five common error types and, accordingly, instantiates five specialized agents—sentiment, rhetoric, dependency, aspect, and reference. Each agent targets the error-prone dimension for LLMs, and their outputs are aggregated and deliberated to produce the final FSA decision.
5.2 Affective Generation
Prompt engineering is also used to improve the performance of AG. The following introduces the application of four prompt engineering technologies in AG.
5.2.1 Zero-shot for Affective Generation
To align prompts with the characteristics of AG, researchers propose Perspective-Taking Lee et al. (2023a), which enhances LLMs’ empathetic dialogue generation by shifting viewpoints. In HEF Yang et al. (2024c), a small empathy model first extracts emotion causes and primary emotion categories. These are then provided to the LLM under zero-shot prompting to produce more precise empathetic responses. The TCG pipeline Bhaskar et al. (2022), built on GPT-3.5-turbo, performs sentence-level opinion extraction, grouping, and clustering to generate long-form summaries of affective opinions.
In the emotion-conditioned text generation task, researchers optimizes prompts by iteratively adding, replacing, and deleting the original prompts Resendiz & Klinger (2023). The optimized prompts better achieve the expected emotions, resulting in improved emotional text generation. Since the output of LLMs largely depends on prompts, users needing more background knowledge might experience a diminished interaction with chatbots due to the absence of contextual understanding. To address this issue, PromptMind Su et al. (2023) generates multiple contextually relevant prompts during the conversation for users to choose from, thereby prompting LLMs to produce answers that better meet user expectations and enhance human-chatbot interaction.
5.2.2 Few-shot for Affective Generation
In the Sibyl framework Wang et al. (2023c), researchers randomly select a single example as a demonstration to prompt the LLM to generate four types of prospective commonsense related to the dialogue, which are then used as training data for subsequent dialogue-generation models. For mixed-initiative dialogue generation tasks such as PersuasionForGood and Emotional Support Conversations, responses are generated conditioned on different exemplar strategies Chen et al. (2023a).
Researchers indicate that generated responses utilizing prompt engineering demonstrate high quality and adherence to semantic controls Chen et al. (2023a). However, in emotional support tasks, the inherent preference of LLMs for specific strategies can hinder effective support despite improving the robustness of predicting appropriate strategies Kang et al. (2024). To address this issue, researchers conduct extensive experiments involving Self-contact approaches, external strategy planning, and example expansion, demonstrating that External-contact approaches help reduce LLM preference bias and enhance the performance of emotional support.
5.2.3 CoT for Affective Generation
In AG, information and knowledge are crucial for producing high-quality responses. To obtain more valuable information, researchers introduce CoT to support empathetic generation. CFEG Chen et al. (2024b) decomposes emotion and reasoning into multiple steps, strengthening the listener’s role awareness and significantly improving the empathy of LLMs. User utterances serve as latent linguistic cues that reveal hidden needs and guide the model to generate more personalized responses. Cue-CoT Wang et al. (2023a) breaks down reasoning into sequential steps to infer user states (e.g., personality, psychological status, emotions) and finally generates responses based on the reasoning outputs. EDIT Wu et al. (2023) enhances response generation by detecting implicit user intentions. It first asks open-ended questions to capture potential intentions, then retrieves knowledge base answers, and finally integrates the additional knowledge into generation. Aptness Hu et al. (2024) adopts a similar strategy by constructing an empathetic response database via multi-step CoT and using retrieval during generation to enhance empathetic capability.
Since AG requires LLMs to address emotion-related issues more proactively, planning information plays an important role in the process of obtaining and using information with LLMs. ProCoT Deng et al. (2023b) guides LLMs to produce intermediate steps of reasoning and planning and analyzes through dynamic reasoning how to achieve dialogue goals, which enhances proactive dialogue ability. ECoT Li et al. (2024g) proposes five consecutive steps to simulate human emotional intelligence and generate responses with richer emotional expression. ESCoT Zhang et al. (2024e) promotes response generation by producing emotional stimuli and designing regulation strategies. Further research transforms step-by-step cognitive commentary prompts into a hierarchical concept Li et al. (2024d) and proposes a three-level emotional integration framework.
Additionally, in metaphor generation, GROUNDS naturally serve as the reasoning chain that connects TENOR and VEHICLE. Researchers guide the model to generate logically consistent GROUNDS Shao et al. (2024a), which in turn produce coherent and rich VEHICLE expressions, thereby improving the quality of metaphor generation.
5.2.4 Agent for Affective Generation
As a new technology based on LLMs, agent is gradually applied to AG. The Agent4SC Zhang et al. (2024f) generates user information, enhances memory, and formulates response strategies through collaboration among LLMs, realistically simulating human sticker usage to improve multimodal empathetic dialogue. The COOPER dialogue framework Cheng et al. (2024a) coordinates multiple specialized agents focusing on specific dialogue goals to generate emotionally supportive and persuasive responses. Researchers propose a multi-turn dialogue framework based on psychologist agents Wu et al. (2024c), in which debaters from different psychological schools generate candidate replies, and an unbiased decision-maker selects the final response, thus addressing the challenge of integrating multiple psychological schools in LLMs. The PerceptiveAgent Yan et al. (2024) simulates different speaking styles according to prompts to build multimodal dialogue systems. The conversational agent with acoustic emotion perception (CA) Hu et al. (2022b) performs speech emotion recognition and incorporates empathetic feedback and interjections to give responses an emotional style. In addition, introducing self-emotion prompts is also shown to help Agents demonstrate more human-like dialogue strategies Zhang et al. (2024c).
Additionally, in the multimodal domain, researchers study multimodal Conversational Health Agents (CHAs) Abbasian et al. (2024) and design five types of agents as components to interact with users to generate audio responses based on user emotions.
5.3 Challenge
Despite the significant achievements of Prompt Engineering in the field of AC, the following challenges persist:
  • As an efficient method with low resource requirements, it greatly reduces the cost of models. However, its performance indicators across various tasks still need to catch up to those of fine-tuned models, unable to meet practical application needs.
  • Currently, the most widely used and advanced CoT method has significantly improved the effectiveness of AC tasks. However, its underlying logic aims to fit human reasoning processes, which conflicts with the unique mechanisms of human emotion generation. This often results in noise during the intermediate reasoning process, affecting the final inference outcomes.
  • Moreover, since emotions are influenced by various factors such as psychology, knowledge, and context, the ways these factors impact and are obtained vary greatly. The current division of agents is not detailed enough, each agent’s design is overly simplistic, and the information aggregation process is quite arbitrary, failing to fully exploit the potential of LLMs.
6 Reinforcement Learning
Reinforcement learning (RL) is increasingly used to align large language models (LLMs) with affect-related objectives beyond supervised fine-tuning. We organize RL for AC into three families that differ by the source of rewards yet share the goal of optimizing behavior at token-, utterance-, and dialogue-levels across AU/AG (see Figure 5): (i) Reinforcement Learning from Human Feedback (RLHF) Ouyang et al. (2022); Rafailov et al. (2023b), where a reward model trained on human comparisons guides policy optimization—often with Proximal Policy Optimization (PPO)—and preference-only variants such as Direct Preference Optimization (DPO) reduce reliance on online RL Ouyang et al. (2022); Schulman et al. (2017); Rafailov et al. (2023b); (ii) Reinforcement Learning with Verifiable/Programmatic Rewards (RLVR) Lambert et al. (2024), which replaces preference labels with rule- or checker-based signals and has inspired algorithms like Group Relative Policy Optimization (GRPO) Shao et al. (2024b) for stable group-wise ranking Lambert et al. (2024); Shao et al. (2024b); and (iii) Reinforcement Learning from AI Feedback (RLAIF), which leverages capable AI evaluators to provide multi-aspect feedback that can be plugged into RL or preference-optimization pipelines. The following subsections detail representative methods and applications of these three families in affective computing.
Figure 5: The overview of reinforcement learning.
6.1 Reinforcement Learning from Human Feedback
A key application of RLHF is enhancing empathy in dialogue systems. Zhu et al. Zhu et al. (2023a) propose a three-stage framework for Empathetic Emotion Elicitation, where reinforcement learning leverages user reactions as rewards to improve emotional engagement across conversations. However, collecting human feedback is costly and difficult to scale. To mitigate this, Xu et al. Xu et al. (2025) introduce Reinforcement Learning from Targeted Human Feedback (RLTHF), which first uses a general-purpose LLM to label data, then directs limited human effort toward “hard” or mislabeled cases. This approach achieves full human-level alignment at only 6–7% of the annotation cost.
The quality of preference data remains critical. Liang et al. Liang et al. (2024) present AlignCap, a preference optimization method that aligns speech emotion captions with human preferences, reducing hallucinations and producing more faithful emotional descriptions. Complementarily, Chen Chen et al. (2025) release EmpathyAgent, the first benchmark for training and evaluating empathetic embodied agents, offering 10,000 multimodal samples with preference-labeled empathetic actions for RLHF training.
6.2 Reinforcement Learning with Verifiable Rewards
To bypass the human annotation bottleneck, RLVR leverages programmatic or model-based reward functions, enabling scalable and automated optimization. A representative work is RL-EMO by Zhang et al. Zhang et al. (2024a), the first to apply RL to multimodal emotion recognition in conversation (ERC). It frames ERC as a sequential decision-making problem, where correct or incorrect emotion predictions yield positive or negative rewards, respectively. Using a Deep Q-Network to model emotion flow, RL-EMO outperforms prior methods and demonstrates the effectiveness of policy optimization for capturing temporal affective dynamics. Zhang et al. Zhang & Jaitly (2025) further propose SAGE (State Augmented GEneration), a framework that explicitly models user and assistant emotional states as latent variables for dialogue generation. Reinforcement learning is applied at the state level, enabling strategic emotional control and producing more empathetic and context-aware responses. Extending into the multimodal domain, Zhao et al. Zhao et al. (2025a) present R1-Omni, which applies RLVR to omni-multimodal (video, audio, text) emotion recognition. With a simple verifiable reward—whether predictions match ground-truth labels—optimized via Group Relative Policy Optimization (GRPO), the framework improves accuracy, robustness, and interpretability in emotion reasoning.
Pushing the boundaries of affective understanding beyond static emotion recognition, the SAGE (Sentient Agent as a Judge) framework Zhang et al. (2025a) reconceptualizes comprehension as a dynamic, interactive process. Its core innovation is a "Sentient Agent" that simulates human-like emotional changes and inner thoughts during multi-turn conversations. By embedding evaluation within the agent’s evolving persona, goals, and emotional state, comprehension is reframed from classification to a stateful simulation of a user "feeling understood." The resulting emotion trajectories and interpretable inner thoughts provide a proxy for modeling and measuring the essence of being heard.
Most RL applications in affective computing focus on generation, steering LLMs toward empathetic, polite, or strategically effective outputs. RLHF naturally fits this domain by embedding human preferences into the training loop. For example, RLVER by Wang et al. Wang et al. (2025b) introduces the first end-to-end RL framework leveraging verifiable emotion rewards from simulated users. Built upon the SAGE simulator, it produces deterministic emotion scores as direct PPO rewards, elevating a 7B open-source model to near-proprietary performance on Sentient-Benchmark. A key innovation in RLVR is finer-grained reward design. Li et al. Li et al. (2024e) propose TOLE, which formulates token-level rewards via classifier probability shifts, enabling efficient controllable text generation. Cai Cai et al. (2024)’s EmpCRL combines empathy and diversity signals, while Li et al. Li et al. (2024c) integrate the "Cognitive Relevance Principle" into ORL, balancing helpfulness against user effort. Together these works show how automated, dense rewards lead to nuanced affective generation. Ma et al. Ma et al. (2025) propose EmpRL, aligning response empathy levels through a programmatic reward from a dedicated "empathy identifier," based on reaction, interpretation, and exploration. Chang et al. Chang et al. (2024) further demonstrate multi-objective reward design in emotional support systems, combining goal-level, flow, and valence rewards for stage-aware dialogue policies. RLVR also supports domain-specific agents. Mishra et al. Mishra et al. (2024) introduce ABLE, a disability support agent guided by a six-component reward model covering persona consistency, politeness, empathy, naturalness, and coherence. Priya et al. Priya et al. (2025) develop GENTEEL-NEGOTIATOR, which uses MoE-based RL with detailed reward functions for polite yet strategic negotiation. For online safety, Wang et al. Wang et al. (2024d) present F2RL, leveraging triplet-based factuality and multi-aspect faithfulness rewards to generate effective counterspeech.
Beyond dialogue, RLVR principles extend into strategic decision-making. Unnikrishnan et al. Unnikrishnan (2024) apply sentiment-aware RL to portfolio management, where sentiment scores from financial news served as programmatic rewards. The resulting trading agent achieved superior profitability, showing how effective signals can optimize complex, non-dialogue tasks.
6.3 Reinforcement Learning from AI Feedback
Within the RLVR paradigm, Reinforcement Learning from AI Feedback (RLAIF) replaces human annotators with capable AI models to provide reward signals. Yoshida et al. Yoshida et al. (2025) exemplify this approach by training an LLM-based reward model on 12 conversational metrics (e.g., empathy, consistency, enjoyability). This AI-generated, multi-faceted feedback is then used with DPO to optimize dialogue agents, showing that AI feedback can scalably capture holistic qualities beyond turn-level scoring.
RLVR also supports new agent architectures. Yuan Yuan et al. (2024) propose ReflectDiffu, integrating RL into diffusion models where a policy network selects empathetic intents conditioned on emotional contagion and mimicry. High-level planning approaches include Deng Deng et al. (2023c)’s Plug-and-Play Policy Planner and Rakib Rakib et al. (2025)’s DialogXpert, which employ LLM critics in self-play to train efficient policy planners for proactive tasks such as emotional support and negotiation. The paradigm further extends to autonomous self-improvement. Ye et al. Ye et al. (2025) propose a self-evolution framework where models refine their own outputs to generate synthetic preference pairs for DPO. Zhao et al. Zhao et al. (2025b) advance this idea with Chain-of-Strategy Optimization (CSO), which uses Monte Carlo Tree Search to explore conversational strategies. Candidate responses are scored by an auxiliary reward model, yielding high-quality preference datasets for DPO fine-tuning, significantly improving strategic empathy. Zhang et al. Zhang et al. (2025b) introduce Decoupled ESC, featuring Inferential Preference Mining (IPM) that identifies psychological errors (e.g., lack of empathy, premature shifts) in model outputs. Erroneous responses are paired with ground-truth ones to build synthetic preference data. The resulting decoupled system, with separate strategy planning and response generation, yields more robust and effective emotional support agents. Collectively, these works illustrate the trajectory of RLAIF toward fully autonomous systems capable of continuously enhancing emotional intelligence.
6.4 Challenge
Despite the growing success of reinforcement learning in aligning LLMs with affective objectives, several key challenges remain:
  • While RLHF has significantly improved model alignment with human preferences, it remains highly resource-intensive, requiring large-scale, high-quality human preference data. This poses a major scalability bottleneck, especially for fine-grained affective tasks that require nuanced and subjective judgments.
  • Many current reward functions used in RLHF and RLVR are designed based on task-level or surface-level attributes (e.g., helpfulness, politeness), but they often lack grounding in psychological or cognitive theories of emotion. This can lead to misalignment between the reward signal and the underlying affective dynamics, limiting the depth of emotional understanding or expression the model can achieve.
  • Despite advances in RL for affective understanding (e.g., RL-EMO, SAGE), temporal affect modeling remains underexplored. Emotions evolve over time, yet most RL-based affective systems still optimize at the turn-level or token-level, failing to capture long-term emotional trajectories in conversation.
  • The lack of standardized benchmarks and evaluation metrics for affective RL makes it difficult to systematically compare approaches or track progress. Current evaluations often rely on subjective human scoring or limited automatic proxies, which may not fully reflect emotional richness, coherence, or alignment with human expectations.
7 Benchmark & Evaluation
This section compiles various benchmarks, evaluation methods, and metrics for Affective Computing, encompassing both Affective Understanding and Affective Generation. Table 2 comprehensively summarizes these benchmarks, detailing the evaluated tasks, datasets, metrics, and models.
Table 2: Various benchmarks for Affective Computing. The AE means Automatic Evaluation, the HE means Human Evaluation.
Name Tasks Datasets Metrics Models
SOUL Deng et al. (2023d) Sentiment Analysis Yelp Zhang et al. (2015), IMDb Maas et al. (2011) AE (F1, BLEU, …), HE Flan-T5, GPT-3.5-turbo
GPT-4V with Emotion Lian et al. (2023b) Sentiment Analysis MVSA-Single, DFEW, etc. 21 datasets AE (Accuracy, WAR, WAF) GPT-4V
MERBench Lian et al. (2024) Multimodal Sentiment Analysis MER2023, etc. 7 datasets AE (WAF, MSE) PLMs, LLMs
DFEW  Jiang et al. (2020) Videol Sentiment Analysis 1500 movies AE(UAR,WAR) EC-STFL
MC-EIU  Liu et al. (2024b) Emotion and Intent Joint Understanding in Multimodal Conversation MC-EIU test AE (WAF) EI2, other PLMs
RLVER Wang et al. (2025b) Empathetic Understanding Sentient Benchmark AE (Sentient Benchmark) GPT-4, GPT-4V, DeepSeek, Claude3.7
SAGE Zhang et al. (2025a) Empathetic Understanding 100-scenario Supportive-Dialogue benchmark AE(BLRI, Sentient emotion score) GPT-4o, Gemini2.5
Cue-CoT Wang et al. (2023a) Emotional Support Conversation ED, etc. 6 datasets AE (Avg BLEU and F1), HE ChatGLM, other LLMs
ESC-EvalZhao et al. (2024a) Emotion Support Conversations 2,801 Role Cards Benchmark AE (ESC-RANK)& HE (7 dimensions) GPT-4, Llama3, EmoLLM, etc.
EmotionBench tse Huang et al. (2023) Empathetic Response Generation EmotionBench test AE, HE GPT series, other LLMs
STICKERCONV Zhang et al. (2024f) Multimodal Empathetic Response Generation STICKERCONV AE(MMr, BLEU, …), HE PEGS, ChatGLM3, Vicuna
SECEU Wang et al. (2023d) Emotional Intelligence Evaluation SECEU test AE (SECEU score, EQ score, …) GPT series, and other LLMs
EQ-Bench Paech (2024) Emotional Intelligence Evaluation EQ-Bench test AE (EQ-Bench scores) GPT series, other LLMs
EmoBench Sabour et al. (2024) Emotional Intelligence Evaluation EmoBench test AE (EmoBench scores) GPT series, other LLMs
EIBENCH Zhao et al. (2024b) Emotional Intelligence, General Intelligence MMLU, etc. 88 datasets AE (Accuracy) Flan-T5, LLaMA-2
EmoVIT Xie et al. (2024) Image Sentiment Explaination Generation EmoVIT test AE (Accuracy) EmoVIT, other MLLMs
EMER Lian et al. (2023a) Explainable Multimodal Emotion Recognition MER2023 AE(Accuracy, BLEU, ROUGE) EMER model, other MLLMs
MER-UniBench Lian et al. (2025) Multimodal Emotion Recognition and Reasoning MER-Caption AE(UAR,WAR), HE AffectGPT, other MLLMs
ChatGPT2AC Amin et al. (2024) Affective Computing res14, et 13 datasets AE(Accuracy, UAR) GPT-4
MM-INSTRUCTEVAL Yang et al. (2024b) Multimodal Reasoning with Multimodal Contexts MVSA-Single etc. 16 datasets AE (Accuracy) GPT series, other LLMs
7.1 Affective Understanding
This section introduces the benchmarks and representative datasets in the field of Affective Understanding for different sentiment analysis-related tasks.
SOUL Deng et al. (2023d) introduces two subtasks: Review Comprehension (RC) to validate subjective statements from reviews, and Justification Generation (JG) to explain sentiment predictions.
GPT-4V with Emotion Lian et al. (2023b) evaluates GPT-4V’s capabilities in Generalized Emotion Recognition (GER), providing quantitative results across 21 benchmark datasets spanning six visual and multimodal emotion recognition tasks.
MERBench Lian et al. (2024) is a benchmark for Multimodal Sentiment Analysis, covering seven prominent datasets. It uses Automatic Evaluation with metrics like weighted average F-score (WAF) and mean squared error (MSE).
DFEW Jiang et al. (2020) is a large-scale dataset for dynamic facial expression recognition, containing over 16,000 video clips from movies with real-world challenges like occlusions and variable poses. Evaluation metrics include UAR and WAR.
MC-EIU Liu et al. (2024b) is a multimodal conversational dataset with nearly 5,000 video clips from bilingual TV series. It targets joint emotion and intent understanding in conversations, evaluated using weighted average F-score (WAF).
RLVER Wang et al. (2025b) is a reinforcement learning framework for enhancing empathetic reasoning in LLMs. It uses the Sentient Benchmark to evaluate a model’s ability to improve a simulated user’s emotional state over time through verifiable emotion rewards.
SAGE Zhang et al. (2025a) is an evaluation framework for higher-order social cognition in LLMs. It uses a simulated Sentient Agent to track emotional changes during dialogues, evaluating a model’s ability to foster positive emotional shifts via its primary metric, the Sentient emotion score.
7.2 Affective Generation
This section introduces the benchmarks in the field of Affective Generation (AG), covering Emotional Support Conversation and Empathetic Response Generation.
7.2.1 Emotional Support Conversation
Emotional support conversation (ESC) aims to alleviate individuals’ emotional intensity and provide guidance for navigating personal challenges through engaging dialogue. Cue-CoT Wang et al. (2023a) is a benchmark consisting of six in-depth dialogue datasets in both Chinese and English, considering three aspects of user statuses: personality, emotions, and psychology. It forms a comprehensive evaluation benchmark for dialogue response generation, employing both Automatic Evaluation and Human Evaluation.
ESC-Eval Zhao et al. (2024a) is a framework for evaluating Emotion Support Conversation (ESC) in LLMs. It uses a specialized role-playing agent (ESC-Role) and a benchmark of 2,801 role cards to simulate multi-turn dialogues. Evaluation is performed via manual annotation across seven dimensions, such as fluency and empathy, enabling a dynamic assessment that moves beyond static, ground-truth-reliant metrics.
7.2.2 Empathetic Response Generation
MEDIC Zhu et al. (2023b) is a multimodal empathy dataset for psychotherapy scenarios, considering visual, audio, and textual modalities. It proposes a conceptual framework for empathy with three mechanisms: Expression of Experience (EE), Emotional Reaction (ER), and Cognitive Reaction (CR).
EmotionBench tse Huang et al. (2023) collects 428 distinct situations categorized into 36 factors and is designed to align LLM emotional responses with human appraisal, focusing on eight negative emotions.
STICKERCONV Zhang et al. (2024f) is a multimodal empathetic dialogue dataset with 2K generated personalities, 12.9K dialogue sessions, and 5.8K unique stickers. It generates textual empathetic responses and outputs stickers as a visual modality. Evaluation is conducted using Automatic Evaluation, Human Evaluation, and LLM-based Evaluation.
7.2.3 Emotional Intelligence
In the context of large language models (LLMs), emotional intelligence (EI) has emerged as a critical area of focus. EI encompasses the capacity to identify, comprehend, regulate, and apply emotions across diverse situations Mayer et al. (2016).
SECEU (Situational Evaluation of Complex Emotional Understanding) Wang et al. (2023d) is a standardized EI test suitable for both humans and LLMs. It evaluates the ability to comprehend and apply emotions in diverse situations.
EQ-Bench Paech (2024) is a benchmark designed to evaluate aspects of EI in LLMs. The test dataset is constructed using GPT-4 to generate dialogues that serve as the context for the test questions.
EmoBench Sabour et al. (2024) is a theory-based comprehensive EI benchmark for LLM evaluation, consisting of 400 hand-crafted questions available in English and Chinese. The framework defines machine EI in terms of emotional understanding and emotional application.
EIBENCH Zhao et al. (2024b) is a large collection of EI-related tasks in a text-to-text format, spanning 15 tasks across 88 datasets. It evaluates three key EI aspects (Emotion Perception, Cognition, and Expression) and general intelligence.
EmoVIT Xie et al. (2024) is the first instruction-tuned benchmark for visual emotion understanding, connecting visual cues with emotional cognition. It features a large-scale dataset of image-instruction pairs for tasks including classification, interpretation, and emotion-grounded conversation.
EMER Lian et al. (2023a) is a benchmark for Explainable Multimodal Emotion Recognition. It requires models to provide explanations for predicted emotions, evaluating their ability to integrate information from audio, video, and text modalities.
MER-UniBench Lian et al. (2025) is the largest benchmark for emotion recognition, with 115K human-verified samples across over 2K fine-grained categories. It supports tasks like fine-grained/basic emotion recognition and sentiment analysis.
7.3 Others
This section introduces other benchmarks that do not easily fit into AU or AG tasks, including affective computing and multimodal reasoning benchmarks.
ChatGPT2AC Amin et al. (2024) evaluates InstructGPT and GPT-4 across a wide range of affective computing tasks, including sentiment analysis, toxicity detection, and personality assessment. The benchmark uses Unweighted Average Recall (UAR) and Accuracy as its automatic evaluation metrics.
MM-INSTRUCTEVAL Yang et al. (2024b) establishes a benchmark for 31 models, including 23 MLLMs, using ten instructions across 16 diverse multimodal reasoning datasets with vision-text contexts. Evaluation metrics include Best Performance, Mean Relative Gain, Stability, and Adaptability.
8 Discussion and Future Works
  1. 1.
    Ethics
    Data is essential and invaluable in artificial intelligence. Sentiment data from the real world, such as data from social media, psychological counseling, is highly personal, sensitive, regional, and cultural. Therefore, ethics must be considered when constructing a dataset, including individual privacy protection, prejudice elimination, and value alignment. Acquiring, preprocessing, and annotating data are all very time-consuming and labor-intensive processes. Thus, leveraging the world knowledge and generative capabilities of LLMs to synthesize datasets for affective computing tasks is an important research direction. Furthermore, improving the quality of original datasets and balancing the new and old datasets is a problem worth studying.
  2. 2.
    Multimodal Affective Computing
    Large Language Models (LLMs), initially designed to work exclusively with text, have now evolved into Multimodal Large Language Models (MLLMs), capable of processing both text and image modalities. In the domain of AC, data from diverse modalities—such as text, visuals, audio, electroencephalogram (EEG), and electrooculogram (EOG)—can convey individual emotions and sentiments. However, most current general-purpose MLLMs face limitations in their ability to integrate and interpret information from multiple modalities in a seamless manner. This often results in misinterpretations, particularly when critical emotional cues are absent or overlooked. To advance AC, it is essential to effectively collect and utilize multimodal data to train unified MLLMs. Key challenges include recognizing micro-expressions, capturing sentiment from audio and video, and enhancing brain-computer interface technologies and emotional intelligence systems. Addressing these challenges represents crucial research directions for the future.
  3. 3.
    Multilingual and Multicultural Affective Computing
    Multilingual and Multicultural sentiment analysis spans different languages and cultures, focusing on sentiment recognition and analysis. Users from different linguistic or cultural backgrounds express emotions differently; for example, the same body movements can have different meanings in various cultural contexts. Current sentiment analysis is typically conducted on a single language or culture. Advances in multicultural sentiment analysis will significantly enhance the understanding of global users’ emotional reactions and behavior patterns, thereby improving the quality of cross-cultural communication and business decision-making. When training the current LLMs, the quality and quantity of corpora in different languages vary greatly, resulting in uneven AC performance in different languages. The core challenge in this field is effectively processing and integrating information from multiple languages and cultures to achieve accurate sentiment analysis. Therefore, it is necessary to continue to improve the multilingual and multicultural capabilities of the LLMs from multiple perspectives.
  4. 4.
    Real-time Affective Computing
    The first step in sentiment analysis is building a dataset and performing sentiment analysis. Real-time sentiment analysis in a dynamic environment, by integrating data in real-time and identifying and analyzing sentiment, is particularly critical for various practical applications such as online customer service, real-time monitoring, and interactive entertainment. Real-time sentiment analysis requires collecting and processing data in real-time, which presents several challenges—ensuring the real-time performance of data processing and sentiment analysis, especially maintaining time synchronization and consistency across different modalities. Real-time systems often encounter various environmental noises. Another major challenge is ensuring that the system can adapt to environmental changes and noise interference while maintaining the accuracy and stability of sentiment recognition.
  5. 5.
    Reasoning in Affective Computing
    Current affective computing tasks primarily focus on understanding emotions and generating emotion-related content. However, they often overlook the deeper, underlying causes that trigger emotional states—specifically, the reasons behind the emergence, persistence, or fluctuation of certain emotions in response to various stimuli or contexts. Gaining insight into these triggers is essential for advancing systems beyond surface-level emotion recognition toward a more nuanced and comprehensive understanding of human affect. A major challenge lies in the extraction and integration of meaningful emotional cues to enable accurate sentiment reasoning. Successfully addressing this challenge could significantly enhance the interpretability and reliability of affective computing systems. However, evaluating and validating the sentiment reasoning remains a significant obstacle, requiring innovative approaches to balance technical accuracy with user comprehension.
  6. 6.
    Effective Evaluation
    Although LLMs have achieved excellent performance in various tasks, it remains essential to explore how to train an LLM that can effectively perceive sentiments and evaluate performance. When training LLMs dedicated to AC tasks, researchers face the challenge of not only retaining the original abilities of models, such as instruction understanding and reasoning, but also strengthening new abilities, specifically sentiment perception and cognition. Additionally, it is crucial to ensure the integrity of values, facts, and logic within the LLMs. In terms of evaluation, previous studies conducted before the emergence of LLMs proposed various metrics to measure model performance, such as Accuracy, BLEU, and ROUGE. However, these metrics are insufficient to fully assess the ability of LLMs, e.g., generalization, generation, perception, and cognitive. Establishing a comprehensive and effective evaluation standard remains an urgent problem to be solved.
9 Conclusion
Affective Computing (AC) has emerged as a crucial research direction in artificial intelligence. With the rise of Large Language Models (LLMs), significant progress has been made in understanding and generating emotions. This paper provides a comprehensive review of LLM applications in AC, exploring the roles of techniques such as instruction tuning, prompt engineering, and reinforcement learning in affective understanding and generation tasks. We further conclude benchmarks for multiple large language models in AC to offer in-depth assistance to researchers and practitioners in related fields. Research indicates that LLMs, leveraging their powerful contextual learning and sequence generation capabilities, have demonstrated excellent affective understanding and generation performance. However, LLMs still face numerous challenges in multilingual and multicultural sentiment analysis, real-time affective computing, and emotion-related tasks in specific vertical domains. We discuss challenges and future research directions related to AC, and aim to provide insights into this rapidly advancing field.
Acknowledgments
Thanks to all co-authors for their hard work. The work is supported by the National Natural Science Foundation of China (No.62272092, No.62172086, and No.62106039), and the Fundamental Research Funds for the Central Universities under Grants (N25XQD004, N25ZLL045).